### v1

# Reason-RFT: Reinforcement Fine-Tuning for Visual Reasoning

Abstract
Visual reasoning abilities play a crucial role in understanding complex multimodal data, advancing both domain-specific applications and artificial general intelligence (AGI). Existing methods improve VLM reasoning via Chain-of-Thought (CoT) supervised fine-tuning, using meticulously annotated training data to enhance visual reasoning capabilities. However, this training paradigm may lead to overfitting and cognitive rigidity, restricting the model’s ability to transfer visual reasoning skills across domains and limiting its real-world applicability. To address these limitations, we propose Reason-RFT, a novel reinforcement fine-tuning framework that significantly enhances generalization capabilities in visual reasoning tasks. Reason-RFT introduces a two-phase training framework for visual reasoning: (1) Supervised Fine-Tuning (SFT) with curated Chain-of-Thought (CoT) data activates the reasoning potential of Vision-Language Models (VLMs), followed by (2) Group Relative Policy Optimization (GRPO)-based reinforcement learning that generates multiple reasoning-response pairs, significantly enhancing generalization in visual reasoning tasks. To evaluate Reason-RFT’s visual reasoning capabilities, we reconstructed a comprehensive dataset spanning visual counting, structure perception, and spatial transformation, serving as a benchmark to systematically assess visual cognition, geometric understanding, and spatial generalization. Experimental results demonstrate Reasoning-RFT’s three key advantages: (1) Performance Enhancement: achieving state-of-the-art results across multiple tasks, outperforming most mainstream open-source and proprietary models; (2) Generalization Superiority: consistently maintaining robust performance across diverse tasks and domains, outperforming alternative training paradigms; (3) Data Efficiency: excelling in few-shot learning scenarios while surpassing full-dataset SFT baselines; Reason-RFT introduces a novel paradigm in visual reasoning, significantly advancing multimodal research. Project website: ReasonRFT.
Figure 1: Overview of Reason-RFT. Compared to traditional SFT-based methods, our proposed Reason-RFT framework demonstrates superior generalization in visual reasoning tasks, excelling in reasoning improvement, out-of-domain performance, and data efficiency.
footnotetext: Equal contribution.footnotetext: Project leaders.footnotetext: ${}^{\text{\Letter}}$ Corresponding author.
1 Introduction
Visual reasoning is pivotal for understanding complex multimodal data and advancing artificial general intelligence (AGI) [38, 45], making it a central focus in intelligent systems research. Recent advancements in image recognition [48, 27] and scene understanding [9, 64] have enabled transformative applications in healthcare [66, 47], robotics [28, 36], and autonomous driving [22, 20, 19, 21]. Consequently, enhancing visual reasoning capabilities has garnered significant attention from both industry and academia for its potential to drive transformative advancements.
Researchers have explored two primary categories of methods to enhance visual reasoning capabilities: (1) neural-symbolic methods [13, 4, 7, 69, 16], which integrate symbolic reasoning with neural networks to improve interpretability and modularity, and (2) Supervised Fine-Tuning (SFT) based on vision-language models (VLMs) [63, 57], which utilize end-to-end training to strengthen reasoning abilities. However, both approaches face significant limitations. Neural-symbolic methods are hindered by high complexity and a strong reliance on program generation, while SFT is constrained by its dependence on high-quality Chain-of-Thought (CoT) annotated data and meticulously designed data mixing strategies, leading to issues such as overfitting, cognitive rigidity, and limited transferability from in-domain to out-of-domain tasks. These challenges reduce their effectiveness in real-world applications.
Recent advances in models such as GPT-o1 [45], DeepSeek-R1 [15], and Kimi-1.5 [56] have demonstrated that large-scale reinforcement learning (RL) during post-training significantly enhances reasoning abilities in mathematics and coding. By providing a dynamic and adaptive alternative to static datasets, RL effectively addresses traditional limitations and improves reasoning, even with minimal data. However, purely RL-based methods face challenges in generalizing across visual reasoning tasks, often limited by the model’s inherent cognitive constraints and the heterogeneity of multimodal data. Motivated by these observations, this paper investigates advanced RL techniques to enhance the adaptability and transferability of visual reasoning, aiming to overcome generalization challenges and improve performance in real-world applications.
In this paper, we propose Reason-RFT, a novel two-phase reinforcement fine-tuning framework designed to enhance generalization in visual reasoning tasks. First, we employ Supervised Fine-Tuning (SFT) with Chain-of-Thought (CoT) reasoning to activate the model’s potential reasoning capabilities, using a high-quality domain-specific visual reasoning dataset tailored to stimulate related reasoning abilities. Subsequently, we further enhance reasoning potential through Group Relative Policy Optimization (GRPO), demonstrating that Reason-RFT achieves superior generalization by pushing the model’s reasoning limits. To evaluate its capabilities, we reconstructed a high-quality dataset spanning visual counting, structure perception, and spatial transformation, serving as a benchmark for assessing visual cognition, geometric understanding, and cross-task generalization. Extensive experiments highlight three key advantages of Reason-RFT: (1) Performance Improvement: It significantly outperforms mainstream VLMs in tasks like visual counting, structure perception, and spatial transformation; (2) Enhanced Generalization: It consistently surpasses SFT-only and RL-only paradigms across diverse tasks, validated through multi-dimensional evaluations; (3) Data Efficiency: It achieves over 95% of the performance of SFT-only approaches using less than 20% of the data. These results underscore the effectiveness and efficiency of Reason-RFT, offering a robust solution for advancing visual reasoning tasks.
Our main contributions are summarized as follows.
  • We propose Reason-RFT, a novel two-phase reinforcement fine-tuning framework that significantly enhances the visual reasoning capabilities of VLMs by effectively combining the complementary strengths of SFT-based and RL-based methods.
  • We systematically analyze SFT-based and RL-based methods in visual reasoning tasks, and highlight the limitations of SFT-based approaches and the superior generalization of RL-based methods in reasoning improvement, out-of-domain performance, and data efficiency.
  • We reconstructed a comprehensive dataset that covers three core domains: visual counting, structure perception, and spatial transformation, which serves as a benchmark to systematically assess visual cognition, geometric understanding, and cross-task generalization.
  • Extensive experiments validate the effectiveness of Reason-RFT, offering valuable insights to advance visual reasoning and introducing a novel paradigm that significantly pushes forward multi-modal research.
2 Related Work
Visual Reasoning Visual reasoning is a key challenge in advancing artificial general intelligence (AGI), requiring models to perform complex cognitive tasks based on visual perception [38, 42, 18]. It has broad applications, including visual counting [38, 37], geometric problem-solving [12, 30, 42, 70, 51], visual transformation reasoning [23], scientific research [41, 31], and robotic task planning [24, 28, 17]. Traditional methods rely on program generation [29, 16, 53] or neural-symbolic approaches [13, 4, 7, 69], while recent vision-language models (VLMs) leverage large language models (LLMs) for enhanced reasoning. For example, LLaVA-CoT [63] uses multi-stage supervised fine-tuning (SFT) with chain-of-thought (CoT) [61], and Insight-V [10] combines SFT with reinforcement learning (RL). DeepSeek-R1-Zero [14] introduces a rule-based RL method, significantly improving reasoning. Building on DeepSeek-R1 [14], our work compares SFT-based and RL-based paradigms, highlighting the superior performance of R1-based methods in visual reasoning.
Post-Training Post-Training is a crucial phase for enhancing the performance of LLMs and VLMs, bridging pre-trained models and their real-world applications [54, 68, 32, 8]. It primarily involves two methodologies: SFT [62, 58] and RL [71, 46, 52, 43, 65]. SFT adapts pre-trained models to specific tasks using task-oriented datasets, often formatted as instructions. Research like FLAN [60] highlights the importance of diverse instruction-tuning datasets for improving zero-shot performance, while iterative processes, such as Llama 3.1’s six-round strategy [11], integrate rejection sampling, synthetic data, and human annotations. RL aligns models with human preferences or task-specific goals through feedback mechanisms. Reinforcement Learning from Human Feedback (RLHF) [46] refines models using human preference data, as seen in Llama 3.1 [11] and Nemotron-4 [2], which use reward modeling techniques like DPO [49] and RPO [2]. For example, TÜLU 3 [34] employs length-normalized DPO, while DeepSeek-V3 [39] combines rule-based and model-based reward systems. Recently, DeepSeek-R1 [14] achieved significant text reasoning improvements through pure RL [50]. Our work adapts R1 methodologies to VLMs, enhancing visual reasoning, and systematically compares SFT-based and R1-based paradigms, demonstrating the superiority of R1-based methods in visual reasoning tasks.
Figure 2: Framework of Reason-RFT. Reason-RFT introduces a two-phase training framework for visual reasoning. First, Supervised Fine-Tuning (SFT) with Chain-of-Thought (CoT) reasoning activates the model’s reasoning potential using a high-quality domain-specific visual reasoning dataset in stage 1. Subsequently, in stage 2, Group Relative Policy Optimization (GRPO) enhances reasoning capabilities further, enabling Reason-RFT to achieve superior generalization by pushing the model’s reasoning limits. Specifically, reward evaluation consists of format reward and three different types of accuracy reward.
3 Methodology
3.1 Problem Definition
Visual reasoning can be formally defined as follows: given a visual input $I$ (e.g., images or videos) and a corresponding textual description or question $T$, the goal is to derive a conclusion or answer $A$ by analyzing the information in the visual input. This process can be represented as a mapping:
$R:(I,T)\rightarrow A,$
where $I\in\mathbb{R}^{H\times W\times C}$ is the visual input with $H$, $W$, and $C$ representing height, width, and channels, $T$ is the textual description or question typically in natural language, and $A$ is the derived conclusion or answer, often in natural language or structured data. Through this mapping, visual reasoning models integrate visual and textual information to achieve effective reasoning.
3.2 Reason-RFT
In this section, we propose Reason-RFT, a novel two-phase hybrid training strategy to enhance the reasoning capabilities of VLMs in complex visual reasoning tasks. As shown in Fig. 2, the framework comprises two stages: (1) SFT-based Visual Reasoning Activation, which uses SFT with high-quality CoT reasoning data to activate the model’s domain-specific reasoning capabilities, and (2) RL-based Visual Reasoning Reinforcement, which employs the GRPO algorithm with rule-based rewards to further push the upper limits of the model’s reasoning potential.
3.2.1 STAGE 1: SFT-based Reasoning Activation
In the initial phase, we employ Supervised Fine-Tuning (SFT) on a structured visual reasoning dataset containing step-by-step reasoning processes. This phase trains the model to decompose complex tasks into logical steps. Each sample is represented as $(x,q,r,a)$, where $x$ denotes the input images, $q$ is the question, $r$ is the reasoning steps, and $a$ is the final answer. The training objective maximizes the likelihood of generating both $r$ and $a$ given $(x,q)$:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,q,r,a)\sim\mathcal{D}}\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}\mid x,q,y_{<t}),$ (1)
where $\mathcal{D}$ is the dataset, $y$ the concatenated sequence of $r$ and $a$, and $\pi_{\theta}$ the model’s token distribution. The output model $\pi_{\text{CoT}}$ serves as the initialization for the next stage, ensuring a robust foundation for reinforcement learning.
3.2.2 STAGE 2: RL-based Reasoning Enhancement
In the second phase, we refine $\pi_{\text{CoT}}$ using Group Relative Policy Optimization (GRPO), leveraging reinforcement learning for its efficiency and scalability. Unlike Proximal Policy Optimization (PPO), which requires a computationally expensive value network, GRPO calculates relative advantages by comparing rewards within a group of sampled actions, reducing computational overhead and simplifying optimization. This makes GRPO particularly suitable for visual reasoning tasks.
Sampling Action Groups For each input state $s=(x,q)$, where $x$ is the visual encoding of the input image and $q$ the textual encoding of the question, GRPO samples a group of actions $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\theta}$, initialized from $\pi_{\text{CoT}}$. The sampling process is:
$a_{i}\sim\pi_{\theta}(a\mid x,q),\quad\text{for }i=1,2,\dots,G.$ (2)
This strategy ensures diverse responses, promoting exploration and preventing premature convergence.
Reward Evaluation Each sampled action $a_{i}$ is assigned a reward $R(a_{i})$ based on verifiable criteria, resulting in a reward set $\{r_{1},r_{2},\dots,r_{G}\}$. For visual reasoning tasks, the reward function $R(a_{i})$ combines two components: Format Reward $R_{\text{format}}(a_{i})$ and Accuracy Reward $R_{\text{acc}}(a_{i})$. The Format Reward ensures structured response formats, while the Accuracy Reward evaluates correctness, balancing structured reasoning and factual accuracy. The reward function is defined as:
$R(a_{i})=R_{\text{format}}(a_{i})+R_{\text{acc}}(a_{i}).$ (3)
Policy Update with Relative Advantage Rewards are normalized within the sampled group to compute relative advantages $\{A_{1},A_{2},\dots,A_{G}\}$, defined as:
$A_{i}=\frac{r_{i}-\text{mean}\{r_{1},r_{2},\dots,r_{G}\}}{\text{std}\{r_{1},r_{2},\dots,r_{G}\}}.$ (4)
Based on these advantages, the policy is updated to reinforce actions with positive advantages and reduce the probability of less effective ones. Policy updates are further constrained by minimizing the KL divergence between the updated and reference models, ensuring stable RL learning.
3.3 Reward Design for Visual Reasoning Tasks
For the diverse requirements of visual reasoning tasks, including Visual Counting, Structure Perception, and Spatial Transformation, our reward design integrates two essential components: Format Reward and Accuracy Reward. The Format Reward is uniformly applied across all tasks, ensuring that the model strictly adheres to a structured response format, which significantly enhances interpretability and consistency. For the Accuracy Reward, we carefully tailor the design to the specific characteristics of each task, as shown in Fig. 2, creating task-specific reward mechanisms to evaluate the correctness of the model’s responses.
Format Reward This component ensures structured and interpretable responses by requiring the model to adhere to a predefined template: reasoning within <think> and </think> and the final answer within <answer> and </answer>. A reward of 1 is given for strict adherence, while deviations result in a reward of 0.
Accuracy Reward This component evaluates the correctness of the model’s responses, ensuring alignment with ground truth across diverse visual reasoning tasks. To address task diversity, we design tailored reward mechanisms for discrete-valued, mathematical, and function-based problems. Each mechanism is crafted to handle the unique characteristics of its problem category, enabling precise and fair evaluation. Below, we introduce the three reward forms.
  • Discrete-valued Type This reward type applies to Visual Counting and partial Structure Perception tasks, where answers are discrete values (e.g., multiple-choice or integer-based responses). The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
    $R_{\text{acc}}(a_{i})=\begin{cases}1,&\text{if }a_{\text{pred}}=a_{\text{gt}}\\ 0,&\text{otherwise},\end{cases}$ (5)
    where $a_{\text{pred}}$ is the model’s predicted answer and $a_{\text{gt}}$ is the ground truth. This binary reward penalizes deviations from the exact ground truth, ensuring precision in tasks requiring unambiguous answers.
  • Mathematical Type This reward type is designed for Structure Perception tasks involving numerical answers, such as floating-point values or LaTeX-formatted expressions. It uses a tolerance-based evaluation to account for minor numerical deviations. The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
    $R_{\text{acc}}(a_{i})=\frac{1}{2}\left[\cos\left(\pi\times\frac{|a_{\text{pred}}-a_{\text{gt}}|-\epsilon_{1}\times|a_{\text{gt}}|}{(\epsilon_{2}-\epsilon_{1})\times|a_{\text{gt}}|}\right)+1\right],$ (6)
    where $a_{\text{pred}}$ is the predicted answer, $a_{\text{gt}}$ is the ground truth, $\epsilon_{1}$ is the tolerance threshold for an exact match (e.g., 0.05), and $\epsilon_{2}$ is the upper bound for partial rewards (e.g., 0.20). If $|a_{\text{pred}}-a_{\text{gt}}|<\epsilon_{1}\times|a_{\text{gt}}|$, the reward is 1 (exact match); if $|a_{\text{pred}}-a_{\text{gt}}|>\epsilon_{2}\times|a_{\text{gt}}|$, the reward is 0 (incorrect). This formulation ensures smooth transitions between full and partial rewards, enabling fair evaluation of numerical accuracy.
  • Function-based Type This reward type is designed for Spatial Transformation tasks requiring a sequence of transformation functions. The accuracy reward $R_{\text{acc}}(a_{i})$ evaluates the alignment between the predicted sequence $T_{\text{pred}}$ and the ground truth $T_{\text{gt}}$, computed as:
    $R_{\text{acc}}(a_{i})=\frac{\text{len}(T_{\text{pred}}^{f+o+v})+\alpha\cdot\text{len}(T_{\text{pred}}^{f+o/v})+\beta\cdot\text{len}(T_{\text{pred}}^{f})}{\max(\text{len}(T_{\text{pred}}),\text{len}(T_{\text{gt}}))},$ (7)
    where $T_{\text{pred}}^{f+o+v}$ is the subset of transformation steps with complete matches (w/ function, object, and value), $T_{\text{pred}}^{f+o/v}$ are the subsets with partial and only-function matches (w/ function and object, or w/ function and value), $T_{\text{pred}}^{f}$ is the subset with only-function matches. $\alpha$ and $\beta$ are the weighting coefficients for partial matches. This formulation ensures nuanced evaluation, rewarding both exact and partially correct responses while allowing flexible adjustment of partial match contributions.
4 Experiment
4.1 Experimental Details
Table 1: Results on three visual reasoning tasks. The best results among different training paradigms are highlighted in bold, while the second-best results are underlined. “ID” denotes in-domain test data, and “OOD” denotes out-of-domain test data.
Method Visual Counting Structure Perception Spatial Transformation
Clevr-Math Super-Clevr AVG GeoMath Geometry3k AVG TRANCE TRANCE-L TRANCE-R AVG
ID OOD ID OOD ID OOD
Proprietary Models
GPT-4o-2024-08-06 [26] 68.10 34.31 51.20 50.18 43.49 46.83 42.55 28.67 29.76 35.88
Gemini-1.5-Pro [55] 61.80 37.50 49.65 50.12 48.38 49.45 26.22 18.76 19.88 22.77
Open-Source Models
Qwen2.5-VL-3B-Instruct [5] 75.90 39.30 57.60 36.75 37.44 37.09 8.57 8.26 8.31 8.42
Phi-3.5-Vision-4B-Instruct [1] 21.40 15.20 18.30 36.83 50.25 43.54 7.42 2.45 4.02 5.33
Llava-OneVision-7B [35] 69.70 29.10 49.40 77.63 43.66 60.64 10.00 8.33 8.74 9.27
Qwen2.5-VL-7B-Instruct [5] 74.60 35.20 54.90 44.00 45.61 44.80 19.63 13.12 13.42 16.45
InternVL-2.5-8B [6] 93.50 35.30 64.40 63.00 47.32 51.60 7.19 6.62 6.63 6.91
Llama-3.2-11B-Vision [44] 10.30 9.50 9.90 13.75 20.85 17.30 8.22 8.40 9.03 8.47
Pixtral-12B [3] 42.60 22.90 32.75 30.38 36.09 33.23 7.35 5.03 5.22 6.42
Qwen2VL-2B-Instruct
Zero-Shot 82.40 32.00 57.20 25.86 20.63 23.25 3.78 4.60 4.67 4.35
$+$ ANS-SFT 96.20 39.20 67.70 51.34 22.50 36.92 77.39 49.24 50.33 58.99
$+$ CoT-SFT 85.50 46.50 66.00 43.05 25.25 34.15 64.37 43.19 42.86 50.14
$+$ Reason-RFT-Zero 98.40 44.80 71.60 47.68 32.50 40.09 42.13 34.07 33.41 33.74
$+$ Reason-RFT 96.80 51.20 74.00 49.03 33.13 41.08 74.61 64.05 64.08 67.58
Qwen2VL-7B-Instruct
Zero-Shot 98.60 42.10 70.35 43.30 43.88 43.59 13.53 12.72 12.78 13.01
$+$ ANS-SFT 95.00 33.90 64.45 51.34 25.38 38.36 82.19 54.29 54.83 63.77
$+$ CoT-SFT 87.30 42.40 64.85 50.49 33.00 41.75 81.31 47.90 47.80 59.00
$+$ Reason-RFT-Zero 99.40 53.00 76.20 55.00 54.75 54.88 67.67 57.20 56.15 56.68
$+$ Reason-RFT 95.60 51.00 73.30 59.27 49.25 54.26 79.97 59.36 58.61 65.98
Datasets In this paper, we comprehensively evaluate the visual reasoning capabilities of our method by leveraging six existing datasets, enhanced through subtask categorization, error-prone data filtering, and dataset restructuring. Detailed protocols for data filtering and restructuring are provided in the supplementary materials. Specifically, we define three task categories as follows.
  • Visual Counting is a multimodal reasoning task evaluating the integration of linguistic, visual, and mathematical skills by solving arithmetic problems in 3D block-based scenes, which includes four subtask types, with datasets from CLEVR-Math [38] and Super-CLEVR [37]. Specifically, CLEVR-Math provides 35k training and 1k test samples, while Super-CLEVR serves as an out-of-domain test set with 1k samples to assess generalization.
  • Structure Perception is a visual reasoning task requiring models to perceive and interpret structural information in mathematical geometries, medical imaging, chart layouts, and architectural designs. It includes both choice and non-choice formats, with datasets sourced from Geo170K [12], Math360K [51] (collectively GeoMath), and Geometry3K [40]. GeoMath provides 4.5k training and 820 test samples, while Geometry3K serves as an out-of-domain test set with 800 samples to evaluate cross-domain adaptability.
  • Spatial Transformation is a spatial-visual reasoning task requiring models to infer single-step or multi-step transformation actions by analyzing initial and final visual states for 3D scenes from multiple perspectives (e.g., center, left, right). The dataset TRANCE [23] features four distinct difficulty levels, with 60k training and 6k test samples. Out-of-domain test samples are generated by rendering identical data from left and right viewpoints to rigorously assess robustness to perspective variations.
Evaluation Metrics We use accuracy-rate (Acc) as the primary metric [67]. For numerical answers, correctness is verified by mathematical equivalence to the ground truth. For multiple-choice questions, we perform a string match. For function-type sequences, we use stepwise multi-level evaluation to assess alignment with the correct solution.
Table 2: More detailed results on Super-Clevr. “Encountered” refers to the types of questions the model has previously seen during Clevr-Math training, while “UnEncountered” denotes the complicated types that the model has not encountered (i.e. questions with mixture of addition and subtraction).
Methods Encountered UnEncountered Total
Qwen2VL-2B-Instruct
Zero-Shot 42.67 0.00 32.00
$+$ ANS-SFT 51.07 5.20 39.20
$+$ CoT-SFT 49.73 36.80 46.50
$+$ Reason-RFT-Zero 58.00 5.20 44.80
$+$ Reason-RFT 60.00 28.40 51.20
Qwen2VL-7B-Instruct
Zero-Shot 54.53 4.80 42.10
$+$ ANS-SFT 42.53 8.00 33.90
$+$ CoT-SFT 45.33 33.60 42.40
$+$ Reason-RFT-Zero 63.60 21.20 53.00
$+$ Reason-RFT 56.13 35.60 51.00
Implementation Details We utilize Qwen2-VL-2B and Qwen2-VL-7B [59] as the backbone models for our experiments. Our implementation is built on the open-source frameworks Open-R1 [25] and vLLM [33], ensuring reproducibility and scalability. All experiments were conducted on a cluster of servers, each equipped with 8$\times$×A800 GPUs. For further details, see the supplementary materials.
Training Paradigms and Baselines To assess the performance and generalization of different training strategies, we compare: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with chain-of-thought (CoT) reasoning; and (2) RL-based methods—Reason-RFT-Zero, which applies reinforcement learning without reasoning activation phase, and Reason-RFT, which uses partial CoT data for reasoning activation before RL training. For comprehensive experiments, we use Qwen2-VL-Instruct [59] as the base model, evaluating both 2B and 7B variants to explore the impact of model scale. In addition, we also select the most advanced open-source models  [5, 1, 35, 6, 44, 3] and the proprietary models  [26, 55] available as baselines to evaluate the performance of different training paradigms.
4.2 Results on In-Domain Tasks
To evaluate the In-Domain (ID) performance of Reason-RFT against different training methods and baseline models across visual reasoning tasks, we conducted extensive training and validation on 2B/7B models for three tasks. The results in Tab. 1 show: (1) Visual Counting RL-based methods outperform all open-source and proprietary baseline models as well as SFT-based methods in both 2B and 7B models, with Reason-RFT-Zero in 7B nodel achieving the best performance; (2) Structure Perception RL-based methods surpass SFT-based approaches in 7B model, while ANS-SFT performs best in 2B model. CoT-SFT shows limited improvement, as enforced reasoning supervision inhibits cognitive enhancement. Additionally, Reason-RFT in 7B model outperforms all proprietary and most open-source models except InternVL-2.5-8B [6] and Llava-OneVision-7B [35]; (3) Spatial Transformation SFT-based methods achieve the highest performance, with Reason-RFT achieving comparable performance while totally surpassing all baseline models. However, Reason-RFT-Zero performs poorly, likely due to its difficulty adapting to function-based tasks, whereas Reason-RFT benefits from reasoning activation, reaching a higher performance ceiling. In summary, Reason-RFT achieves performance comparable to SFT-based methods in ID tasks while outperforming all baselines in Visual Counting and Spatial Transformation, and most of baselines in Structure Perception.
4.3 Results on Out-of-Domain Generalization
To validate the Out-of-Domain (OOD) performance of Reason-RFT across different training methods and baseline models in visual reasoning tasks, we conducted comprehensive experiments on 2B/7B models for three tasks. Results in Tab. 1 and Tab. 2 show that: (1) Visual Counting RL-based methods demonstrate superior generalization over SFT-based methods in both 2B and 7B models, with Reason-RFT surpassing ANS-SFT by 12% (2B) and 17% (7B) while outperforming all open-source and proprietary baselines. Notably, compared with purely RL-based method (i.e., Reason-RFT-Zero), CoT-SFT and Reason-RFT can activate model capabilities on unseen complicated questions, achieving strong results despite its poor zero-shot performance before in-domain training, as shown in Tab. 2. This also highlights the limitations of Reason-RFT-Zero due to cognitive constraints, especially in smaller models (i.e. 2B models); (2) Structure Perception RL-based methods consistently outperform SFT-based methods, with Reason-RFT achieving the best result in 2B model (8% higher than CoT-SFT) while Reason-RFT-Zero performs strongest in 7B model and Reason-RFT remains competitive (16% higher than CoT-SFT), with SFT-based methods showing limited impact especially in the 7B model; (3) Spatial Transformation RL-based methods surpass SFT-based approaches in both 2B and 7B models while significantly outperforming all baseline models, with Reason-RFT (2B) demonstrating remarkable OOD generalization exceeding GPT-4o [26] by 35%. In summary, Reason-RFT surpasses all open-source and proprietary baselines as well as alternative training methods, demonstrating remarkable performance in visual reasoning generalization capabilities.
4.4 Assessment on Training Efficiency
To demonstrate the data efficiency of Reason-RFT during training, we trained all methods on TRANCE and recorded intermediate and validation results, as shown in Fig. 3. Detailed results are provided in the Appendix. Key findings include: (1) Reason-RFT exhibits strong data efficiency. In the ID tasks of the 2B model, Reason-RFT achieves 70% of the performance of Reason-RFT-Zero using only 3% of the training data (1,600 samples), reaching 82.5% with 9% of the data. (2) This efficiency generalizes to OOD tasks. In the 7B model, Reason-RFT achieves over 92% of the performance of Reason-RFT-Zero using just 3% of the training data, demonstrating robust generalization capabilities.
Figure 3: Results of different methods on the Spatial Transformation task across training processes. (a) Evaluation results for 2B model on ID task, (b) Evaluation results for 7B model on ID task, (c) Evaluation results for 2B model on OOD task, (d) Evaluation results for 7B model on OOD task.
4.5 Training Insights
Transient Adaptation Gap This phenomenon refers to the temporary performance drop observed during the initial phase of training process with Reason-RFT-Zero. As the model transitions from directly outputting answers to generating structured reasoning processes, it experiences a brief adaptation difficulty, leading to a sharp performance decline followed by gradual recovery. As shown in Fig. 4 (a), which illustrates the training process of Reason-RFT-Zero on the CLEVR-MATH, both ID and OOD test performances exhibit this sharp drop and recovery within the first 100 steps. We further investigate this phenomenon through a case study, shown in Fig. 4 (b), which demonstrates that forcing the model to output its reasoning process can lead to incorrect answers to originally correct question.
Greedy Reward Stratification This phenomenon describes the model’s tendency during Reason-RFT-Zero training to prioritize easier rewards (e.g., Format Reward) over harder ones (e.g., Accuracy Reward). As shown in Fig. 5, the model’s Reasoning Token Length initially decreases, then gradually increases before stabilizing. This behavior coincides with the Format Reward reaching its initial peak and the Accuracy Reward entering its rapid growth phase. We infer that the model simplifies its outputs early on to quickly adapt to the structured response format, reducing Reasoning Token Length. Once the Format Reward is maximized, the model shifts focus to improving accuracy, increasing Reasoning Token Length.
Figure 4: Illustration of the Transient Adaptation Gap. (a) shows a sharp drop and recovery in both ID and OOD test performances within the early training steps during training on the CLEVR-MATH task. (b) presents a case study of the prediction result on the 10-th step for further investigation.
Figure 5: Analysis of Greedy Reward Stratification. The model’s Reasoning Token Length initially decreases, then gradually increases before stabilizing during Reason-RFT-Zero training. (a) illustrates the curve of Reasoning Token Length, highlighting the Minimum Point. (b) shows the curve of Format Reward, indicating the Maximum Point. (c) displays the curve of Accuracy Reward, marking the Rapid Growth Point. This trend aligns with the Format Reward reaching its initial peak and the Accuracy Reward entering its rapid growth phase.
Figure 6: Analysis of Reasoning Redundancy. (a) illustrates the curves of Reasoning Token Length for Reason-RFT-Zero and Reason-RFT during training on the Structure Perception task, highlighting the significant difference in Reasoning Token Length between these two paradigms in the later steps of training. (b) shows the curves of Accuracy Reward, indicating that the accuracy rates of both paradigms converge and remain nearly identical.
Reasoning Redundancy This phenomenon refers to the significant difference in Reasoning Token Length between models trained under different paradigms. For example, in the Structure Perception task, Reason-RFT and Reason-RFT-Zero achieve comparable final test accuracy, but Reason-RFT exhibits notably longer Reasoning Token Length, as shown in Fig. 6. This occurs because the CoT data used in Reason-RFT’s Reasoning Activation phase, often distilled from stronger models (e.g., GPT-4o), leads the model to learn longer reasoning chains during CoT-SFT. During the Reasoning Reinforcement phase, the model retains these lengthy chains due to the lack of penalties or incentives for response length. In contrast, Reason-RFT-Zero, which lacks CoT data, stabilizes at a shorter Reasoning Token Length through exploration. We hypothesize that Reason-RFT’s longer reasoning chains may be unnecessary for the current task difficulty, introducing redundant computational overhead. Experiments limiting Reasoning Token Length during inference for Reason-RFT-trained models show stable performance until a certain threshold, partially confirming Reasoning Redundancy in specific tasks.
4.6 Exploration on Reward Design
Table 3: Results of different format reward strategies on the Spatial Transformation task.
Setting ID OOD AVG
TRANCE TRANCE-L TRANCE-R
Qwen2VL-2B-Instruct
Reason-RFT-Zero 42.13 34.07 33.41 33.74
$+$ visual tokens 42.01 36.05 35.97 38.01
Reason-RFT 74.61 64.05 64.08 69.33
$+$ visual tokens 71.99 60.13 59.87 65.99
Qwen2VL-7B-Instruct
Reason-RFT-Zero 67.67 57.2 56.15 62.17
$+$ visual tokens 70.28 59.52 57.01 64.27
Reason-RFT 79.97 59.36 58.61 69.48
$+$ visual tokens 79.85 58.71 57.98 69.09
Table 4: Results of different accuracy reward strategies on the Spatial Transformation task.
Setting $\alpha$ $\beta$ ID OOD AVG
TRANCE TRANCE-L TRANCE-R
Qwen2VL-2B-Instruct
Baseline 0 0 74.61 64.05 64.08 69.33
(a) 0.50 0.25 79.18 56.36 55.45 67.54
(b) -0.25 -0.50 73.69 64.41 64.72 69.13
Qwen2VL-7B-Instruct
Baseline 0 0 79.97 59.36 58.61 69.48
(a) 0.50 0.25 80.89 53.20 52.61 66.90
(b) -0.25 -0.50 75.03 64.83 63.18 69.52
Table 5: Results of various mixed CoT activation datasets of Reason-RFT on the Structure Perception task.
Setting CoT Activation Data ID OOD AVG
GeoMath Geometry3k
Baseline GeoMath-only data 59.27 49.25 54.26
(a) Mixed Specific-Domain data 50.61 45.35 48.02
(b) Mixed General-Domain data 42.51 40.25 41.38
Exploration on Format Reward In DeepSeek-R1 [15], the Format Reward mandates the use of <think> tags for structuring reasoning processes in pure-textual tasks. To enhance visual reasoning tasks, we propose integrating visual observations through captioning, extending the format reward with <summary> and <caption> tokens alongside <think>. As shown in Tab. 3, this extension significantly improved the performance of Reason-RFT-Zero, but not Reason-RFT. We attribute this discrepancy to Reason-RFT’s prior exposure to Reasoning Activation, which likely learned the caption-like guidance from a limited CoT reasoning samples, resulting in reduced effectiveness of additional tag guidance. Conversely, Reason-RFT-Zero, being purely RL-based, benefits more from effective tag guidance, highlighting its greater potential for improvement.
Exploration on Accuracy Reward We investigate the Accuracy Reward for the challenging Spatial Transformation task, which requires predicting novel transformation sequences. The reward design in Eq. 7 includes coefficients $\alpha$ and $\beta$ for partial matches, influencing tolerance for incomplete outputs. We test three configurations: (1) $\alpha=0,\beta=0$ (full matches only), (2) $\alpha=0.50,\beta=0.25$ (full and partial matches), and (3) $\alpha=-0.25,\beta=-0.50$ (full matches with partial penalties). Experimental results on 2B and 7B models (Tab. 4) revealed: (1) Rewards for partial matches improved in-domain performance but harmed out-of-domain generalization, suggesting “soft rewards” may limit adaptability; (2) Penalties for partial matches enhanced performance on out-of-domain, indicating “hard rewards” are more effective for serialized tasks.
4.7 Exploration on COT Activation Data
To investigate the impact of differently composed CoT activation data on Reason-RFT, we construct two distinct datasets: a mixed domain-specific dataset, which integrates relevant yet distinct data from in-domain tasks, and a mixed general-domain dataset, encompassing a broader range of visual reasoning tasks (e.g., graph topology, visual puzzles). Using these datasets, we perform Reason-RFT training on Structure Perception task, with the results detailed in Tab. 5. From this, two key points emerge: (1) As the proportion of in-domain training data decreases, the model’s performance on specific tasks declines; (2) Models trained on more diverse visual reasoning domain data may also exhibit a reduction in domain-specific performance.
5 Conclusion
In this paper, we propose Reason-RFT, a novel reinforcement fine-tuning framework that enhances the generalization capabilities of visual reasoning models. By integrating supervised fine-tuning (SFT) with Chain-of-Thought (CoT) reasoning activation data and Group Relative Policy Optimization (GRPO)-based reinforcement learning, Reason-RFT effectively mitigates key challenges such as overfitting and cognitive rigidity, thereby improving cross-domain transferability and real-world applicability. To support systematic evaluation, we reconstruct a comprehensive dataset covering visual counting, structure perception, and spatial transformation tasks, establishing a robust benchmark for assessing model performance across diverse scenarios. Extensive experiments demonstrate the effectiveness of Reason-RFT, providing valuable insights for advancing visual reasoning research and introducing a new paradigm in multimodal learning.
Appendix
This supplementary material provides additional details on the proposed method and experimental results that could not be included in the main manuscript due to page limitations. Specifically, this appendix is organized as follows.
  • Sec. A provides more details on the evaluation of reasoning tasks and discusses how we collected, filtered, and reconstructed a high-quality dataset.
  • Sec. B outlines the models and training processes, providing more detailed experimental specifics.
  • Sec. C presents comprehensive experimental results.
  • Sec. D includes more visualization cases.
Appendix A Details of Evaluation Reasoning Tasks
A.1 Visual Counting
Task Definition Visual Counting is a multi-modal reasoning task that evaluates the integration of language, visual, and mathematical capabilities by solving basic arithmetic problems in visual scenes featuring 3D blocks with diverse attributes like color, size, material, and shape. The task includes four different types: 1) Subtraction Counting after removing specific object subsets, 2) Addition Counting after inserting objects with specified quantities and attributes, 3) Adversarial Designed as trick questions where actions are performed on one object, but the query targets an unaffected object, and 4) Multi-Hop Counting after sequential steps by addition or subtraction action. This task challenges models to perform attribute-based reasoning in dynamic visual contexts, emphasizing cross-modal understanding and reasoning capabilities. Some examples are shown in Fig. 7.
Figure 7: The sample of Visual Counting.
Dataset Preparation For in-domain dataset, we refined the original dataset from CLEVR-Math [38] by filtering out low-quality or incorrect samples using GPT-4o, resulting in a clean dataset comprising 35K training samples and 1K test samples. These samples are categorized into four specific types: subtraction, addition, adversarial, and multihop-subtraction. To evaluate out-of-domain generalization, we extended CLEVR-Math by enhancing the diversity of objects through the incorporation of 3D assets from Super-CLEVR [37], which leads to the creation of Super-CLEVR-Math, an advanced benchmark with 1K test samples designed to assess model generalization under increased complexity. These test samples are also divided into four task types: addition, subtraction, addition-subtraction, and subtraction-multihop. Notably, the addition-subtraction type introduces a novel category consisting both addition and subtraction, which is not present in CLEVR-Math, further elevating the benchmark’s challenge and diversity.
Reward Design Following the reward methodology of DeepSeek-R1 [15], we define two distinct reward functions: Format Reward and Accuracy Reward. The Format Reward is assigned a value of 1 if the response adheres to the predefined template structure, specifically in the form of <think>...</think><answer>...</answer>; otherwise, it is assigned a value of 0. The Accuracy Reward is assigned a value of 1 if the numerical counting result in the response is correct; otherwise, it is assigned a value of 0. This dual-reward mechanism ensures both structural compliance and numerical accuracy in model responses.
A.2 Structure Perception
Task Definition Structure Perception represents a complex class of visual mathematical reasoning tasks, which focuses on assessing the model’s capacity to determine geometric structure relationships and perform calculations involving angles, lengths, areas, and other geometric properties. The task includes problems such as identifying congruent or similar shapes, calculating perimeters and areas, determining angles between lines or shapes, and solving problems related to geometric transformations (e.g., rotations, translations, and reflections). By combining mathematical rigor with visual reasoning, this task challenges models to demonstrate a deep understanding of geometric principles in both abstract and real-world scenarios. Some examples are shown in Fig. 8.
Figure 8: The sample of Structure Perception.
Dataset Preparation For the in-domain dataset, we utilized GeoMath-8K, a dataset specifically designed for geometric problem-solving, which is constructed based on Math360K [51] and Geo170K [12]. To ensure data quality, we employed GPT-4o to filter out incorrect samples and removed those with answers that were neither numerical nor included in the provided options, thereby streamlining the validation process during training and testing. This refinement process resulted in a curated dataset consisting of 4.5K training samples and 820 test samples. For out-of-domain evaluation, we selected 800 samples from Geometry3K [40] (including 400 multiple-choice and 400 open-ended questions) to comprehensively assess the model’s generalization capabilities on geometry reasoning.
Reward Design We maintain the same Format Reward as used in the Visual Counting task above. The Accuracy Reward is extended to support the evaluation of both multiple-choice questions and mathematical expressions, ensuring comprehensive assessment across various problem types. Specifically, mathematical reward type is designed for Structure Perception tasks involving numerical answers, such as floating-point values or LaTeX-formatted expressions. It uses a tolerance-based evaluation to account for minor numerical deviations. The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
$R_{\text{acc}}(a_{i})=\frac{1}{2}\left[\cos\left(\pi\times\frac{|a_{\text{pred}}-a_{\text{gt}}|-\epsilon_{1}\times|a_{\text{gt}}|}{(\epsilon_{2}-\epsilon_{1})\times|a_{\text{gt}}|}\right)+1\right],$ (8)
where $a_{\text{pred}}$ is the predicted answer, $a_{\text{gt}}$ is the ground truth, $\epsilon_{1}$ is the tolerance threshold for an exact match (e.g., 0.05), and $\epsilon_{2}$ is the upper bound for partial rewards (e.g., 0.20). If $|a_{\text{pred}}-a_{\text{gt}}|<\epsilon_{1}\times|a_{\text{gt}}|$, the reward is 1 (exact match); if $|a_{\text{pred}}-a_{\text{gt}}|>\epsilon_{2}\times|a_{\text{gt}}|$, the reward is 0 (incorrect). This formulation ensures smooth transitions between full and partial rewards, enabling fair numerical evaluation.
A.3 Spatial Transformation
Task Definition Spatial Transformation is a spatial-visual reasoning task designed to infer single-step or multi-step transformation actions by analyzing the initial and final visual states from multiple perspectives (e.g., center, left, right). The task utilizes transformation functions, including change_size, change_color, change_material, change_shape, and change_position, to modify object properties such as size, color, material, shape, and position using predefined values. This task evaluates the model’s ability to reason about spatial relationships and object transformations across diverse viewpoints in dynamic visual scenarios. Some examples are shown in Fig. 9.
Dataset Preparation We generated 100K samples using the environment and configuration from Trance [23], with each sample comprising initial object attributes, front-view image of initial state, and images of final state captured from front, left, and right perspectives. To ensure high data quality, we implemented a rigorous filtering process: (1) removing samples containing occluded or invisible objects in either the initial or final states, (2) eliminating redundant actions within the transformation sequences, and (3) consolidating multi-step displacement actions, which collectively ensure the uniqueness and correctness of the solutions. The refined dataset consists of 60K training samples and 6K test samples. For the training set, we constructed the Trans-Center-60K dataset using the Center-Center configuration, which pairs front-view initial and final state images. For in-domain evaluation, we derived the Trans-Center-6K dataset from the 6K test samples under the same Center-Center configuration. To evaluate out-of-domain generalization, we constructed two additional datasets: Trans-Left-6K and Trans-Right-6K, leveraging the Center-Left and Center-Right configurations to systematically assess the model’s generalization capabilities in spatial reasoning under diverse viewpoint conditions.
Figure 9: The sample of Spatial Transformation.
Reward Design For the Format Reward, we adopted the same formulation as used in the Visual Counting task. As for the Accuracy Reward, a specialized design was developed to evaluate the sequence of transformation functions. Function-based type is designed for Spatial Transformation tasks requiring a sequence of transformation functions. The accuracy reward $R_{\text{acc}}(a_{i})$ evaluates the alignment between the predicted sequence $T_{\text{pred}}$ and the ground truth $T_{\text{gt}}$, computed as:
$R_{\text{acc}}(a_{i})=\frac{\text{len}(T_{\text{pred}}^{f+o+v})+\alpha\cdot\text{len}(T_{\text{pred}}^{f+o/v})+\beta\cdot\text{len}(T_{\text{pred}}^{f})}{\max(\text{len}(T_{\text{pred}}),\text{len}(T_{\text{gt}}))},$ (9)
where $T_{\text{pred}}^{f+o+v}$ is the subset of transformation steps with complete matches (w/ function, object, and value), $T_{\text{pred}}^{f+o/v}$ are the subsets with partial and only-function matches (w/ function and object, or w/ function and value), $T_{\text{pred}}^{f}$ is the subset with only-function matches. $\alpha$ and $\beta$ are the weighting coefficients for partial matches. This formulation ensures nuanced evaluation, rewarding both exact and partially correct responses while allowing flexible adjustment of partial match contributions.
System Prompts For the Spatial Transformation task, we designed two versions of the system prompt. The first version specifies the answer output format using the <think> and <answer> tags, while the second version includes additional outputs <summary> and <caption> for experiments on exploration of format reward design in main paper. These two versions are illustrated in Fig. 10 and Fig. 11, respectively.
Appendix B Details of Models and Training
We utilize Qwen2-VL-2B and Qwen2-VL-7B [59] as the backbone models for our experiments. Our implementation is built on the open-source frameworks Open-R1 [25] and vLLM [33], ensuring reproducibility and scalability. All experiments were conducted on a cluster of servers, each equipped with 8$\times$×A800 GPUs. For the Visual Counting task and Spatial Transformation task, we trained the models for 1 epoch each on their respective training datasets, ensuring sufficient exposure to task-specific patterns while avoiding overfitting. For the Structure Perception task, due to its GeoMath training dataset consisting of a relatively small number of training samples (a total of 4,500), we extended the training duration to 6 epochs to allow the models to fully capture the underlying structural and geometric relationships. In the Reason-RFT training pipeline, all models underwent an initial CoT activation stage with 1,600 samples before proceeding to the RL phase.
Appendix C More Experiment Results
Performance at Different Training Samples Fig. 12 and Fig. 13 illustrate the In-Domain (ID) and Out-of-Domain (OOD) performance of all methods across three visual reasoning tasks, evaluated at various training sample sizes. This analysis helps us understand how each method scales with training data. More detail evaluation results for each subset of three tasks are in Tab. 6 - Tab. 13.
We systematically varied the number of training samples, from minimal to substantial, allowing us to identify performance thresholds and data efficiency for each method in both ID and OOD contexts. Key findings from this analysis include: Data Efficiency of Reason-RFT: Reason-RFT demonstrates exceptional data efficiency, achieving approximately 70% of the performance of Reason-RFT-Zero with only 3% of the training data (1,600 samples), and 82.5% with just 9Robust Generalization to OOD Tasks: In the 7B model, Reason-RFT achieves over 92% of Reason-RFT-Zero’s performance using just 3% of the training data, showcasing its strong generalization capabilities. Comparison Across Methods: Reason-RFT consistently outperforms other methods, particularly in data-constrained scenarios, indicating its suitability for applications with limited data availability. Performance Saturation: As training sample size increases, some methods experience performance plateaus, suggesting that beyond a certain point, additional data yields diminishing returns.
In conclusion, the evaluation of performance across different training samples not only highlights the strengths of Reason-RFT in terms of data efficiency and generalization but also provides critical insights into the performance dynamics of various methods. These findings are essential for practitioners aiming to maximize performance while effectively managing training resources.
Appendix D Visualization
In this section, we present additional visualization results, see Fig. 14 - Fig. 22. Reason-RFT demonstrates superior performance over CoT-SFT in terms of logical consistency, reasoning quality, and correctness. CoT-SFT’s flaws stem from incorrect assumptions and misinterpretations, highlighting the importance of accurate problem interpretation and reasoning in visual reasoning tasks.
Figure 10: The system prompt used in Spatial Transformation task.
Figure 11: The system prompt used in Spatial Transformation task w/ <summary> and <caption> tags in format.
Figure 12: Results of all methods on Qwen2VL-2B-Instruct, ID and OOD performance at different training checkpoints.
Figure 13: Results of all methods on Qwen2VL-7B-Instruct, ID and OOD performance at different training checkpoints.
Methods Steps Visual Counting
Clevr-Math (ID)
adversarial sub-multi addition subtraction AVG
Zero-Shot - 93.60 84.00 55.60 96.40 82.40
ANS-SFT 100 83.60 56.40 91.20 81.60 78.20
200 69.20 67.60 91.60 82.00 77.60
400 81.60 65.60 90.80 84.80 80.70
600 72.40 73.20 92.40 89.20 81.80
800 78.40 77.20 82.80 90.40 82.20
1200 85.60 78.00 91.60 95.60 87.70
1600 92.80 82.40 94.80 96.80 91.70
2187 95.20 92.80 97.60 99.20 96.20
CoT-SFT 100 49.20 40.00 82.00 69.20 60.10
200 65.20 55.60 88.00 76.40 71.30
400 66.00 57.20 90.00 79.60 73.20
600 67.20 59.20 87.20 82.80 74.10
800 77.60 61.60 92.40 85.20 79.20
1200 76.80 70.00 91.20 93.60 82.90
1600 80.80 66.80 91.60 92.00 82.80
2187 83.20 71.20 93.20 94.40 85.50
Reason-RFT-Zero 100 92.80 88.80 94.40 96.00 93.00
200 95.60 91.60 95.60 97.60 95.10
400 92.00 87.60 84.00 96.40 90.00
600 94.40 92.80 93.60 96.00 94.20
800 96.40 96.40 96.00 98.80 96.90
1200 98.40 95.60 100.00 99.60 98.40
1600 96.40 94.80 98.80 99.60 97.40
2500 98.40 95.60 99.60 100.00 98.40
Reason-RFT 100 89.60 73.20 93.60 95.60 88.00
200 89.20 78.00 95.20 96.40 89.70
400 92.80 82.40 95.20 97.60 92.00
600 94.80 86.00 96.80 97.20 93.70
800 96.80 88.40 96.80 98.80 95.20
1200 94.80 86.00 96.40 98.80 94.00
1600 94.40 91.60 97.20 99.60 95.70
2500 98.40 92.80 96.80 99.20 96.80
Table 6: Complete experimental results of Qwen2VL-2B-Instruct on the Clevr-Math test set after training on Clevr-Math. “sub-multi” donates the subtraction-multihop task.
Methods Steps Visual Counting
Super-Clevr (OOD)
addition subtraction add-sub sub-multi AVG Encountered UnEncountered
Zero-Shot - 10.40 54.40 0.00 63.20 32.00 42.67 0.00
ANS-SFT 100 51.20 37.60 11.60 39.20 34.90 42.67 11.60
200 38.40 55.60 8.40 15.60 29.50 36.53 8.40
400 40.80 45.20 5.60 35.20 31.70 40.40 5.60
600 41.20 61.60 8.00 35.60 36.60 46.13 8.00
800 49.20 50.40 7.20 26.00 33.20 41.87 7.20
1200 44.00 53.20 5.60 38.80 35.40 45.33 5.60
1600 48.80 53.60 6.00 26.00 33.60 42.80 6.00
2187 49.60 62.00 5.20 41.60 39.60 51.07 5.20
CoT-SFT 100 47.20 50.00 28.80 25.60 37.90 40.93 28.80
200 56.00 52.40 38.00 34.00 45.10 47.47 38.00
400 55.20 57.20 22.40 30.40 41.30 47.60 22.40
600 58.40 55.20 24.00 35.60 43.30 49.73 24.00
800 57.60 47.60 26.80 41.60 43.40 48.93 26.80
1200 58.00 54.40 35.60 32.40 45.10 48.27 35.60
1600 53.20 58.40 33.20 40.40 46.30 50.67 33.20
2187 53.60 58.80 36.80 36.80 46.50 49.73 36.80
Reason-RFT-Zero 100 46.00 65.20 6.80 58.80 44.20 56.67 6.80
200 48.80 66.00 9.20 57.60 45.40 57.47 9.20
400 42.00 71.20 8.40 50.80 43.10 54.67 8.40
600 47.20 65.20 7.60 47.60 41.90 53.33 7.60
800 56.40 69.20 6.80 55.20 46.90 60.27 6.80
1200 52.00 73.60 7.20 59.20 48.00 61.60 7.20
1600 51.60 71.60 6.40 54.80 46.10 59.33 6.40
2500 49.60 71.20 5.20 53.20 44.80 58.00 5.20
Reason-RFT 100 59.20 57.60 38.00 41.60 49.10 52.80 38.00
200 59.60 64.40 39.20 42.00 51.30 55.33 39.20
400 61.60 64.00 39.20 37.20 50.50 54.27 39.20
600 66.80 67.20 32.00 46.00 53.00 60.00 32.00
800 66.00 65.60 34.00 39.20 51.20 56.93 34.00
1200 67.20 65.20 33.60 40.80 51.70 57.73 33.60
1600 63.60 66.00 33.20 44.80 51.90 58.13 33.20
2500 68.00 67.20 28.40 44.80 52.10 60.00 28.40
Table 7: Complete experimental results of Qwen2VL-2B-Instruct on the Super-Clevr test set after training on Clevr-Math. “add-sub” donates the addition-subtraction task, while “sub-multi” donates the subtraction-multihop task.
Methods Steps Visual Counting
Clevr-Math (ID)
adversarial sub-multi addition subtraction AVG
Zero-Shot - 99.60 98.40 97.60 98.80 98.60
ANS-SFT 100 69.20 54.00 81.20 69.20 68.40
200 61.20 50.00 82.40 75.60 67.30
400 69.20 63.60 89.20 77.60 74.90
600 70.40 54.00 90.40 81.20 74.00
800 80.00 74.00 91.20 89.20 83.60
1200 86.80 79.20 94.40 91.20 87.90
1600 90.40 84.40 95.20 92.00 90.50
2187 96.80 89.20 96.80 97.20 95.00
CoT-SFT 100 81.60 63.60 91.20 83.60 80.00
200 80.00 64.00 92.00 88.80 81.20
400 72.40 66.00 88.80 79.60 76.70
600 77.60 66.00 94.40 89.20 81.80
800 78.40 65.20 94.00 87.20 81.20
1200 79.60 76.80 92.40 88.00 84.20
1600 86.40 78.00 92.80 93.20 87.60
2187 87.20 78.80 93.60 89.60 87.30
Reason-RFT-Zero 100 98.00 94.40 98.80 99.60 97.70
200 99.60 93.20 99.20 100.00 98.00
400 99.60 95.20 99.60 98.80 98.30
600 98.00 98.40 100.00 99.60 99.00
800 99.60 98.40 99.60 98.80 99.10
1200 100.00 98.00 99.60 99.20 99.20
1600 99.60 97.60 100.00 99.20 99.10
2500 99.60 98.40 100.00 99.60 99.40
Reason-RFT 100 88.80 79.20 95.60 94.40 89.50
200 92.00 80.00 96.40 95.20 90.90
400 94.40 84.40 96.00 95.60 92.60
600 92.80 84.00 96.40 97.60 92.70
800 92.80 85.20 96.80 96.40 92.80
1200 94.80 89.60 97.20 97.60 94.80
1600 94.80 86.40 97.60 97.20 94.00
2500 96.80 88.40 99.20 98.00 95.60
Table 8: Complete experimental results of Qwen2VL-7B-Instruct on the Clevr-Math test set after training on Clevr-Math. “sub-multi” donates the subtraction-multihop task.
Methods Steps Visual Counting
Super-Clevr (OOD)
addition subtraction add-sub sub-multi AVG Encountered UnEncountered
Zero-Shot - 46.80 75.20 4.80 41.60 42.10 54.53 4.80
ANS-SFT 100 57.60 41.20 5.60 46.40 37.70 48.40 5.60
200 42.00 38.80 8.00 33.60 30.60 38.13 8.00
400 37.20 46.40 5.20 31.60 30.10 38.40 5.20
600 32.00 44.80 12.40 19.20 27.10 32.00 12.40
800 38.80 38.00 6.80 37.20 30.20 38.00 6.80
1200 42.00 42.80 12.80 32.00 32.40 38.93 12.80
1600 36.40 48.40 11.20 17.20 28.30 34.00 11.20
2187 39.60 58.80 8.00 29.20 33.90 42.53 8.00
CoT-SFT 100 60.00 63.60 44.00 41.60 52.30 55.07 44.00
200 67.60 66.40 48.00 46.80 57.20 60.27 48.00
400 55.20 60.40 19.60 42.00 44.30 52.53 19.60
600 64.80 61.20 35.20 43.20 51.10 56.40 35.20
800 60.00 53.60 37.60 42.40 48.40 52.00 37.60
1200 51.20 56.00 35.20 39.60 45.50 48.93 35.20
1600 53.20 56.40 34.40 35.20 44.80 48.27 34.40
2187 51.60 51.60 33.60 32.80 42.40 45.33 33.60
Reason-RFT-Zero 100 58.80 82.80 24.00 62.40 57.00 68.00 24.00
200 56.00 83.20 18.80 50.00 52.00 63.07 18.80
400 62.40 79.60 22.80 37.60 50.60 59.87 22.80
600 61.20 85.20 17.20 49.20 53.20 65.20 17.20
800 52.80 86.80 20.40 52.00 53.00 63.87 20.40
1200 53.60 83.20 19.20 46.80 50.70 61.20 19.20
1600 61.20 84.80 18.40 43.20 51.90 63.07 18.40
2500 59.20 86.40 21.20 45.20 53.00 63.60 21.20
Reason-RFT 100 53.60 56.80 33.20 39.60 45.80 50.00 33.20
200 52.00 61.20 31.60 44.00 47.20 52.40 31.60
400 56.00 59.60 30.80 45.20 47.90 53.60 30.80
600 56.00 64.00 31.60 50.00 50.40 56.67 31.60
800 56.00 60.00 28.00 41.60 46.40 52.53 28.00
1200 66.00 65.60 38.00 50.40 55.00 60.67 38.00
1600 64.40 59.60 32.40 48.80 51.30 57.60 32.40
2500 62.80 60.80 35.60 44.80 51.00 56.13 35.60
Table 9: Complete experimental results of Qwen2VL-7B-Instruct on the Super-Clevr test set after training on Clevr-Math. “add-sub” donates the addition-subtraction task, while “sub-multi” donates the subtraction-multihop task.
Methods Steps Structure Perception
Geometry3k (OOD) GeoMath (ID)
CHOICE NON-CHOICE AVG CHOICE NON-CHOICE AVG
Zero-Shot - 40.25 1.00 20.63 35.57 20.31 25.86
ANS-SFT 100 35.25 16.25 25.75 58.72 29.89 40.37
200 33.25 17.50 25.38 56.38 35.44 43.05
400 30.75 17.00 23.88 64.77 35.06 45.86
600 - - - 73.83 38.12 51.10
800 32.75 16.00 24.38 72.15 36.40 49.39
1200 - - - 73.83 35.44 49.39
1600 29.00 16.00 22.50 74.83 37.36 50.98
1686 28.75 16.25 22.50 74.83 37.93 51.34
CoT-SFT 100 16.50 21.50 19.00 31.54 34.10 33.17
200 7.50 23.50 15.50 32.89 35.25 34.39
400 21.50 21.25 21.38 41.61 40.04 40.61
600 - - - 43.62 36.59 39.14
800 16.50 23.50 20.00 45.97 39.27 41.70
1200 - - - 53.02 40.04 44.76
1600 24.25 24.00 24.13 53.69 37.93 43.66
1686 26.75 23.75 25.25 51.34 38.31 43.05
Reason-RFT-Zero 100 32.25 17.75 25.00 41.61 31.23 35.00
200 33.00 18.50 25.75 48.99 35.06 40.12
400 41.50 23.50 32.50 52.68 34.87 41.34
600 37.00 22.75 29.88 60.74 37.55 45.98
800 42.25 25.00 33.63 62.42 40.42 48.42
1200 43.00 23.75 33.38 61.07 39.66 47.44
1600 42.75 22.25 32.50 63.09 38.31 47.32
1610 43.25 21.75 32.50 63.09 38.89 47.68
Reason-RFT 100 37.50 23.25 30.38 50.34 41.00 44.39
200 33.50 29.25 31.38 56.71 40.04 46.10
400 38.25 28.75 33.50 56.38 39.27 45.49
600 40.50 27.25 33.88 61.41 41.19 48.54
800 41.25 29.50 35.38 58.05 41.19 47.32
1200 40.25 31.00 35.63 61.74 42.34 49.39
1600 38.00 29.25 33.63 62.08 43.10 50.00
1610 36.75 29.50 33.13 60.74 42.34 49.03
Table 10: Complete experimental results of Qwen2VL-2B-Instruct on the Structure Perception task after training on GeoMath.
Methods Steps Structure Perception
Geometry3k (OOD) GeoMath (ID)
CHOICE NON-CHOICE AVG CHOICE NON-CHOICE AVG
Zero-Shot - 45.25 23.00 34.13 61.07 38.12 46.46
ANS-SFT 100 38.50 18.25 28.38 64.77 34.87 45.74
200 32.50 22.75 27.63 69.46 35.25 47.68
400 - - - 72.48 40.42 52.07
600 32.25 18.00 25.13 73.49 39.27 51.71
800 - - - 75.50 37.93 51.58
1200 32.50 18.50 25.50 75.84 37.74 51.59
1600 32.50 18.25 25.38 75.84 37.36 51.34
1686 18.25 38.75 28.50 38.59 42.72 41.22
CoT-SFT 100 6.50 32.00 19.25 38.26 43.10 41.34
200 27.00 34.50 30.75 56.71 44.64 49.03
400 - - - 52.68 44.06 47.19
600 35.50 36.25 35.88 63.09 43.49 50.61
800 - - - 63.42 42.91 50.36
1200 29.50 37.50 33.50 64.09 44.06 51.34
1600 29.25 36.75 33.00 61.74 44.06 50.49
1686 58.50 41.75 50.13 56.71 45.98 49.88
Reason-RFT-Zero 100 59.00 44.25 51.63 63.42 45.21 51.83
200 62.00 43.00 52.50 70.47 45.40 54.51
400 - - - 70.13 46.74 55.24
600 64.75 45.25 55.00 70.47 49.23 56.95
800 - - - 66.11 46.17 53.42
1200 69.00 43.25 56.13 71.14 45.59 54.88
1600 66.25 43.25 54.75 69.80 46.55 55.00
1610 46.75 37.50 42.13 67.79 45.79 53.79
Reason-RFT 100 53.00 37.00 45.00 72.82 46.93 56.34
200 52.75 37.25 45.00 71.14 46.55 55.49
400 51.50 37.00 44.25 73.49 48.28 57.44
600 56.75 37.25 47.00 77.52 46.17 57.56
800 59.00 40.00 49.50 79.87 48.08 59.63
1200 56.00 39.50 47.75 74.50 49.62 58.66
1600 59.00 39.50 49.25 78.52 48.28 59.27
1610 59.00 39.50 49.25 78.52 48.28 59.27
Table 11: Complete experimental results of Qwen2VL-7B-Instruct on the Structure Perception task after training on GeoMath.
Method Steps Spatial Transformation
TRANCE (ID) TRANCE-L (OOD) TRANCE-R (OOD)
Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG
GPT-4o / 47.28 42.96 40.87 39.08 42.55 23.16 30.56 30.73 30.22 28.67 24.38 31.74 31.13 31.77 29.76
Zero-Shot / 2.10 3.27 4.08 5.68 3.78 2.02 4.73 5.57 6.08 4.60 2.39 4.73 5.59 5.95 4.67
ANS-SFT 100 15.90 19.12 14.67 13.12 15.70 10.60 11.33 10.01 9.07 10.25 11.08 12.17 10.17 10.07 10.87
200 23.97 29.56 33.98 33.95 30.37 13.25 26.76 31.26 32.58 25.96 13.53 26.54 30.53 31.20 25.45
400 44.95 42.58 40.75 33.65 40.48 26.03 35.98 34.26 29.38 31.41 24.06 35.62 35.46 31.73 31.72
800 62.10 56.55 53.01 47.55 54.80 24.70 42.27 43.05 42.65 38.17 26.09 38.98 42.95 42.55 37.64
1600 80.70 75.68 68.34 64.60 72.33 34.38 49.00 53.83 53.10 47.58 35.84 49.28 51.23 52.50 47.21
3200 82.85 80.30 78.00 71.60 78.19 36.22 52.61 55.23 54.27 49.58 38.51 52.78 54.47 53.98 49.94
final 82.70 79.93 76.70 70.22 77.39 36.00 52.82 54.59 53.55 49.24 39.63 53.75 54.33 53.60 50.33
COT-SFT 100 6.99 14.90 15.99 20.36 14.56 10.32 13.74 11.82 14.69 12.64 6.97 13.38 12.84 13.41 11.65
200 15.45 19.12 14.53 16.46 16.39 12.90 17.51 14.22 15.82 15.11 11.23 17.71 13.87 16.32 14.78
400 25.98 26.74 19.94 16.02 22.17 15.73 21.19 17.73 15.96 17.65 16.13 20.55 17.91 15.97 17.64
800 43.85 43.19 41.77 37.84 41.66 22.88 38.29 37.54 35.29 33.50 22.74 35.81 37.60 36.23 33.10
1600 52.82 61.06 54.38 45.85 53.53 28.65 43.95 40.99 40.13 38.43 29.00 41.23 40.78 39.26 37.57
3200 61.40 69.15 65.32 62.28 64.54 28.67 45.97 50.06 52.10 44.20 31.19 45.87 45.92 51.35 43.58
final 67.47 67.52 62.78 59.70 64.37 28.87 44.41 49.16 50.30 43.19 30.20 44.77 47.15 49.33 42.86
Reason-RFT-Zero 100 8.44 17.96 20.69 26.22 18.33 8.53 17.42 21.16 25.05 18.04 8.08 18.12 21.09 25.70 18.25
200 9.59 18.76 22.97 28.73 20.01 9.49 20.08 23.19 27.00 19.94 9.72 18.97 22.82 28.50 20.00
400 12.35 21.47 27.01 26.25 21.77 11.10 21.47 25.73 25.30 20.90 10.54 21.19 25.44 25.60 20.69
800 18.47 32.08 32.77 27.85 27.79 15.40 29.12 30.93 27.38 25.71 15.52 27.75 31.50 27.88 25.66
1600 36.78 40.20 37.78 34.51 37.32 19.96 33.03 35.84 34.49 30.83 20.39 32.85 33.87 33.90 30.25
3200 43.72 46.89 44.07 40.50 43.80 18.67 34.11 37.85 39.69 32.58 18.08 34.01 37.27 40.27 32.41
final 46.21 45.01 44.53 42.11 44.47 18.33 34.57 37.94 40.57 32.85 18.28 33.45 37.44 40.34 32.38
Reason-RFT 100 53.52 55.47 58.91 53.35 55.31 31.84 47.02 50.62 50.39 44.97 31.29 46.08 48.61 49.85 43.96
200 54.97 59.77 63.67 59.46 59.47 35.72 49.28 54.85 54.67 48.63 36.74 52.27 53.68 54.16 49.21
400 63.80 66.97 68.47 64.70 65.99 39.74 55.94 61.27 57.94 53.72 41.10 56.35 59.16 57.22 53.46
800 64.33 68.13 66.88 63.15 65.62 47.40 61.64 63.00 58.45 57.62 46.76 60.60 61.40 59.91 57.17
1600 76.47 73.42 74.05 69.16 73.28 52.68 62.22 66.56 64.71 61.54 53.47 65.17 64.98 63.52 61.79
3200 72.88 74.85 75.77 72.45 73.99 52.58 63.60 68.51 66.01 62.68 52.67 65.60 67.60 65.66 62.88
final 74.10 74.52 76.68 73.12 74.61 53.49 65.72 69.64 67.34 64.05 54.95 66.25 68.32 66.80 64.08
Table 12: Complete experimental results of Qwen2VL-2B-Instruct on the Spatial Transformation task after training on TRANCE.
Method Steps Spatial Transformation
TRANCE (ID) TRANCE-L (OOD) TRANCE-R (OOD)
Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG
GPT-4o / 47.28 42.96 40.87 39.08 42.55 23.16 30.56 30.73 30.22 28.67 24.38 31.74 31.13 31.77 29.76
Zero-Shot / 16.25 16.42 10.96 10.48 13.53 11.71 16.80 11.50 10.85 12.72 13.30 16.08 10.55 11.18 12.78
ANS-SFT 100 40.30 37.05 30.67 28.35 34.09 32.07 31.12 26.00 26.80 29.00 26.38 29.71 27.74 26.48 27.58
200 65.18 53.33 49.43 45.15 53.27 33.29 45.14 45.61 45.52 42.39 35.17 43.43 45.97 43.02 41.90
400 65.33 59.35 57.17 50.77 58.16 32.40 44.13 47.69 46.23 42.61 32.10 44.88 47.04 45.25 42.32
800 78.90 70.67 63.97 62.10 68.91 34.08 50.62 51.99 52.95 47.41 34.22 50.40 50.62 52.88 47.03
1600 78.50 76.12 73.80 66.25 73.67 38.85 52.97 57.93 56.05 51.45 37.77 53.57 56.45 55.92 50.93
3200 83.80 83.23 82.83 78.17 82.01 40.10 56.02 61.02 59.90 54.26 40.78 55.06 61.67 60.98 54.62
final 83.70 84.10 82.50 78.45 82.19 39.67 55.58 61.84 60.05 54.29 42.64 54.84 61.44 60.38 54.83
COT-SFT 100 20.58 28.98 25.97 30.00 26.38 21.89 29.94 29.18 29.33 27.59 18.49 28.43 30.43 30.36 26.93
200 41.80 44.08 46.02 42.16 43.52 25.31 36.28 36.81 38.86 34.32 21.19 34.55 37.06 37.54 32.59
400 45.39 51.32 58.20 52.42 51.83 32.53 44.61 47.73 46.10 42.74 31.87 38.96 44.62 45.90 40.34
800 54.87 61.97 62.20 59.93 59.74 30.19 46.01 50.01 52.88 44.77 29.91 45.02 49.60 52.83 44.34
1600 71.27 71.14 72.82 69.93 71.29 28.82 46.43 51.01 58.94 46.30 29.08 45.25 52.24 58.31 46.22
3200 84.13 80.62 79.99 78.42 80.79 29.93 47.63 56.49 62.54 49.15 30.46 47.85 54.83 61.02 48.54
final 86.50 79.43 80.54 78.77 81.31 28.07 47.54 54.42 61.58 47.90 29.69 45.32 54.69 61.48 47.80
Reason-RFT-Zero 100 23.59 31.62 33.22 31.27 29.93 15.88 26.86 28.13 30.19 25.27 15.21 27.29 27.54 29.88 24.98
200 35.06 39.45 36.80 34.77 36.52 20.39 30.22 31.15 31.20 28.24 18.10 29.27 30.57 30.81 27.19
400 25.28 40.78 41.70 35.35 35.78 20.20 39.28 35.43 33.44 32.09 21.72 39.47 37.63 32.79 32.90
800 50.18 51.55 50.43 46.06 49.56 35.44 46.15 45.65 39.55 41.70 33.90 45.89 46.48 41.57 41.96
1600 59.60 61.90 57.30 55.36 58.54 43.95 55.03 52.96 50.60 50.64 41.28 56.46 51.64 49.08 49.62
3200 62.50 68.53 68.79 66.22 66.51 42.54 58.05 58.97 60.10 54.92 42.56 56.93 60.02 60.21 54.93
final 65.63 68.30 69.45 67.30 67.67 46.61 58.22 61.69 62.26 57.20 45.53 58.40 61.81 58.85 56.15
Reason-RFT 100 60.07 62.95 68.53 62.78 63.58 37.09 52.54 60.21 54.99 51.21 35.26 52.56 57.30 55.42 50.14
200 67.69 66.98 69.53 66.00 67.55 36.26 53.94 60.20 60.05 52.61 34.33 54.21 58.25 60.79 51.90
400 74.72 71.31 73.62 69.14 72.20 36.11 56.00 62.25 63.01 54.34 36.95 55.72 60.82 64.88 54.59
800 72.57 71.04 75.58 73.26 73.11 38.93 56.86 64.10 63.69 55.90 35.06 54.39 60.82 64.88 53.79
1600 79.18 76.70 76.59 73.26 76.43 37.25 53.44 66.38 66.76 55.96 35.01 55.39 66.02 66.67 55.77
3200 79.70 79.22 80.13 76.61 78.92 37.63 55.73 65.24 68.02 56.66 37.18 56.34 63.90 68.89 56.58
final 80.48 79.15 81.97 78.26 79.97 43.68 61.04 64.66 68.05 59.36 45.45 56.34 63.33 69.33 58.61
Table 13: Complete experimental results of Qwen2VL-7B-Instruct on the Spatial Transformation task after training on TRANCE.
Figure 14: Visualization examples of Reason-RFT and CoT-SFT on Visual Counting task.
Figure 15: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 16: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 17: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 18: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 19: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 20: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 21: Visualization examples of Reason-RFT and CoT-SFT on Spatial Transformation task.
Figure 22: Visualization examples of Reason-RFT and CoT-SFT on Spatial Transformation task.