### v3

# Reason-RFT: Reinforcement Fine-Tuning for Visual Reasoning of Vision Language Models

Abstract
Visual reasoning abilities play a crucial role in understanding complex multimodal data, advancing both domain-specific applications and artificial general intelligence (AGI). Existing methods improve Vision-Language Models (VLMs) reasoning via Chain-of-Thought (CoT) supervised fine-tuning, using meticulously annotated training data to enhance visual reasoning capabilities. However, this training paradigm may lead to overfitting and cognitive rigidity, restricting the model’s generalization ability to transfer visual reasoning skills under domain shift and limiting its real-world applicability. To address these limitations, we propose Reason-RFT, the first two-stage reinforcement fine-tuning framework for visual reasoning: (1) Supervised Fine-Tuning (SFT) with curated CoT data activates the reasoning potential of VLMs, followed by (2) Group Relative Policy Optimization (GRPO)-based reinforcement learning that generates multiple reasoning-response pairs, significantly enhancing the capability to address ubiquitous domain shift in visual reasoning tasks. To evaluate the visual reasoning capabilities of Reason-RFT, we reconstructed a comprehensive dataset encompassing visual counting, structural perception, and spatial transformation, serving as a benchmark for systematic assessment across three core dimensions. Experimental results demonstrate three key advantages: (1) Performance Enhancement: achieving state-of-the-art results across multiple tasks, outperforming mainstream open-source and proprietary models; (2) Generalization Superiority: consistently maintaining robust performance in addressing domain shift in typical visual reasoning tasks, outperforming alternative paradigms; (3) Data Efficiency: excelling in few-shot learning scenarios while surpassing full-dataset SFT baselines. Reason-RFT introduces a rebust training paradigm in visual reasoning, and please refer to project website: Reason-RFT.
footnotetext: Equal contribution. Project leader. $^{\text{\Letter}}$ Corresponding author: shanghang@pku.edu.cn.
1 Introduction
Visual reasoning is pivotal for understanding complex multimodal data and advancing artificial general intelligence (AGI) [1, 2], making it a central focus in intelligent systems research. Recent advancements in image recognition [3, 4, 5, 6, 7, 8, 9], interactive security [10, 11, 12] and scene understanding [13, 14] have enabled transformative applications in healthcare [15, 16], robotics [17, 18, 19, 20, 21, 22, 23, 24], and autonomous driving [25, 26, 27, 28, 29, 30, 31]. Consequently, enhancing visual reasoning capabilities has garnered significant attention from both industry and academia for its potential to drive transformative advancements.
Researchers have explored two primary categories of methods to enhance visual reasoning capabilities: (1) neural-symbolic methods [32, 33, 34, 35, 36], which integrate symbolic reasoning with neural networks to improve interpretability and modularity, and (2) Supervised Fine-Tuning (SFT) based on vision-language models (VLMs) [37, 38], which utilize end-to-end training to strengthen reasoning abilities. However, both approaches face significant limitations. Neural-symbolic methods are hindered by high complexity and a strong reliance on program generation, while SFT is constrained by its dependence on high-quality Chain-of-Thought (CoT) annotated data and meticulously designed data mixing strategies, leading to issues such as overfitting, cognitive rigidity, and limited adaptability to domain shift. These challenges reduce their effectiveness in real-world applications.
Recent advances such as GPT-o1 [2], DeepSeek-R1 [39], and Kimi-1.5 [40] show that reinforcement learning (RL) during post-training enhances reasoning in coding and mathematics. RL offers a dynamic alternative to SFT by enabling exploration and feedback-driven optimization, which can improve performance with limited labeled data. However, pure RL methods often lack robustness to domain shifts—such as changes in visual appearance or configuration, limiting their generalization capacity in real-world visual reasoning scenarios.
Figure 1: Overview of Reason-RFT. Compared to traditional SFT-based methods, our proposed Reason-RFT framework demonstrates superior generalization in visual reasoning tasks, excelling in reasoning improvement, domain shift adaptability, and data efficiency.
To address this, we propose Reason-RFT, the first two-stage reinforcement fine-tuning framework designed to enhance generalization in visual reasoning tasks. First, we employ SFT with CoT reasoning to activate the model’s potential reasoning capabilities, using a high-quality domain-specific visual reasoning dataset tailored to stimulate related reasoning abilities. Subsequently, we further enhance reasoning potential through Group Relative Policy Optimization (GRPO), demonstrating that Reason-RFT improves robustness under distribution shifts by enhancing the model’s reasoning capabilities. To evaluate its effectiveness, we constructed a high-quality dataset covering visual counting, structure perception, and spatial transformation, serving as a benchmark for evaluating three core capabilities of visual reasoning. Extensive experiments highlight three key advantages of Reason-RFT: (1) Performance Improvement: It achieves strong results on visual reasoning tasks, including visual counting, structure perception, and spatial transformation, outperforming mainstream VLMs; (2) Enhanced Generalization: It consistently exceeds both SFT-only and RL-only baselines under domain shift conditions, as demonstrated through comprehensive evaluations; (3) Data Efficiency: It reaches over 90% of the SFT-only performance while using less than 5% of the data. These results underscore the effectiveness and efficiency of Reason-RFT, establishing it as a robust framework for advancing visual reasoning. Our main contributions are summarized as follows.
  • We introduce Reason-RFT, a two-stage reinforcement fine-tuning framework that significantly enhances the visual reasoning capabilities of VLMs by effectively combining the complementary strengths of SFT-based and RL-based methods.
  • We provide a systematic analysis of SFT-based and RL-based paradigms on visual reasoning tasks, identifying the limitations of SFT and the advantages of RL in improving reasoning ability, handling domain shifts, and achieving data-efficient learning.
  • We reconstruct a comprehensive dataset spanning three core domains: visual counting, structure perception, and spatial transformation, serving as a benchmark for evaluating visual cognition, geometric understanding, and spatial generalization.
  • Extensive experiments validate the proposed framework, demonstrating its practicality and effectiveness, and providing a new perspective for reinforcement-driven multi-modal training.
2 Related Work
Visual Reasoning Visual reasoning is a core challenge in advancing AGI, requiring models to perform complex cognitive tasks grounded in visual perception [1, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It underpins a wide range of applications, including visual counting [1, 51], geometric problem-solving [52, 53, 41, 54, 55], visual transformation reasoning [56, 57], scientific analysis [58, 59], and robotic task planning [60, 17, 61]. Traditional approaches rely on program generation [62, 36, 63] or neural-symbolic reasoning [32, 33, 34, 35], while recent advances in VLMs leverage large language models (LLMs) to enhance reasoning capabilities. For instance, LLaVA-CoT [37] employs multi-stage SFT with CoT prompting [64], and Insight-V [65] integrates SFT with RL. DeepSeek-R1-Zero [66] further introduces a rule-guided RL framework that substantially improves reasoning performance. Building upon the DeepSeek-R1 [66], our work provides a comparative analysis of SFT-based and RL-based paradigms, demonstrating the advantages of R1-style methods in enhancing visual reasoning.
Post-Training Post-Training is a crucial phase for enhancing the performance of LLMs and VLMs, bridging pre-trained models and their real-world applications [67, 68, 69, 70]. It primarily involves two methodologies: SFT [71, 72] and RL [73, 74, 75, 76, 77]. SFT adapts pre-trained models to specific tasks using task-oriented datasets, often formatted as instructions. Research like FLAN [78] highlights the importance of diverse instruction-tuning datasets for improving zero-shot performance, while iterative processes, such as Llama 3.1’s six-round strategy [79], integrate rejection sampling, synthetic data, and human annotations. RL aligns models with human preferences or task-specific goals through feedback mechanisms. Reinforcement Learning from Human Feedback (RLHF) [74] refines models using human preference data, as seen in Llama 3.1 [79] and Nemotron-4 [80], which use reward modeling techniques like DPO [81] and RPO [80]. For example, TÜLU3 [82] employs length-normalized DPO, while DeepSeek-V3 [83] combines rule-based and model-based reward systems. Recently, DeepSeek-R1 [66] achieved significant text reasoning improvements through pure RL [84]. Our work first adapts R1 methodologies to VLMs, enhancing visual reasoning, and systematically compares SFT-based and RL-based paradigms in visual reasoning tasks.
3 Methodology
In this section, we introduce Reason-RFT, a novel two-stage training strategy to enhance the reasoning capabilities of VLMs in complex visual reasoning tasks. As shown in Fig. 2, the framework comprises two stages: (1) SFT-based Visual Reasoning Activation, which uses SFT with high-quality CoT reasoning data to activate the model’s domain-specific reasoning capabilities, and (2) RL-based Reasoning Enhancement, which employs the GRPO algorithm with rule-based rewards to further push the upper limits of the model’s reasoning potential.
3.1 STAGE 1: SFT-based Reasoning Activation
In the initial stage, we employ SFT on a structured visual reasoning dataset containing step-by-step reasoning processes. This stage trains the model to decompose complex tasks into logical steps. Each sample is represented as $(x,q,r,a)$, where $x$ denotes the input images, $q$ is the question, $r$ is the reasoning steps, and $a$ is the final answer. The training objective maximizes the likelihood of generating both $r$ and $a$ given $(x,q)$:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,q,r,a)\sim\mathcal{D}}\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}\mid x,q,y_{<t}),$ (1)
where $\mathcal{D}$ denotes the dataset, $y$ represents the concatenated sequence of $r$ and $a$, and $\pi_{\theta}$ denotes the model’s token distribution. The resulting model $\pi_{\text{CoT}}$ is used to initialize the subsequent stage, providing a stable foundation for RL-based reasoning enhancement.
3.2 STAGE 2: RL-based Reasoning Enhancement
In the second stage, we refine $\pi_{\text{CoT}}$ using GRPO, leveraging RL for its efficiency and scalability. Unlike Proximal Policy Optimization (PPO), which requires a computationally expensive value network, GRPO calculates relative advantages by comparing rewards within a group of sampled actions, reducing computational overhead and simplifying optimization. This makes GRPO particularly suitable for visual reasoning tasks.
Sampling Action Groups For each input state $s=(x,q)$, GRPO samples a group of actions $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\theta}$, initialized from $\pi_{\text{CoT}}$. The sampling process is:
$a_{i}\sim\pi_{\theta}(a\mid x,q),\quad\text{for }i=1,2,\dots,G.$ (2)
This strategy ensures diverse responses, promoting exploration and preventing premature convergence.
Reward Evaluation. Each sampled action $a_{i}$ receives a reward $R(a_{i})$ based on verifiable criteria, yielding a reward set $\{r_{1},r_{2},\dots,r_{G}\}$. For visual reasoning tasks, the reward $R(a_{i})$ is composed of a format reward $R_{\text{format}}(a_{i})$, which enforces structured outputs, and an accuracy reward $R_{\text{acc}}(a_{i})$, which measures correctness. This formulation balances structural alignment and factual precision in reasoning. The reward function is defined as:
$R(a_{i})=R_{\text{format}}(a_{i})+R_{\text{acc}}(a_{i}).$ (3)
Policy Update with Relative Advantage Rewards are normalized within the sampled group to compute relative advantages $\{A_{1},A_{2},\dots,A_{G}\}$, defined as:
$A_{i}=\frac{r_{i}-\text{mean}\{r_{1},r_{2},\dots,r_{G}\}}{\text{std}\{r_{1},r_{2},\dots,r_{G}\}}.$ (4)
Based on these advantages, the policy is updated to reinforce actions with positive advantages and reduce the probability of less effective ones. To maintain training stability, the update is constrained by minimizing the KL divergence between the updated and reference policies.
Figure 2: Framework of Reason-RFT. Reason-RFT adopts a two-stage training paradigm for visual reasoning. The first stage applies SFT with CoT reasoning to establish strong task-specific priors. In the second stage, GRPO is employed to further enhance reasoning capability and generalization.
3.3 Reward Design for Visual Reasoning Tasks
For the diverse requirements of visual reasoning tasks, including visual counting, structure perception, and spatial transformation, our reward design integrates two essential components: format reward and accuracy reward. The format reward is uniformly applied across all tasks, ensuring that the model strictly adheres to a structured response format for consistency. For the accuracy reward, we carefully tailor the design to the specific characteristics of each task, as shown in Fig. 2, creating task-specific reward mechanisms to evaluate the correctness of the model’s responses.
Format Reward This component ensures structured and interpretable responses by requiring the model to adhere to a predefined template: reasoning within <think> and </think> and the final answer within <answer> and </answer>. A reward of 1 is only given for strict adherence.
Accuracy Reward This component evaluates the correctness of the model’s responses, ensuring alignment with ground truth across diverse visual reasoning tasks. To address task diversity, we design tailored reward mechanisms for discrete-valued, mathematical, and function-based problems. Each mechanism is crafted to handle the unique characteristics of its problem category, enabling precise and fair evaluation. Below, we introduce the three reward forms.
  • Discrete-valued Type This reward type applies to visual counting and structure perception tasks, where answers are discrete values (e.g., multiple-choice or integer-based responses). The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
    $R_{\text{acc}}(a_{i})=\begin{cases}1,&\text{if }a_{\text{pred}}=a_{\text{gt}}\\ 0,&\text{otherwise},\end{cases}$ (5)
    where $a_{\text{pred}}$ is the predicted answer and $a_{\text{gt}}$ is the ground truth. This discrete reward penalizes deviations from ground truth, ensuring precision in tasks requiring unambiguous answers.
  • Mathematical Type This reward type is designed for structure perception tasks involving numerical answers, such as floating-point values or LaTeX-formatted expressions. It uses a tolerance-based evaluation to account for minor numerical deviations, which is defined as:
    $R_{\text{acc}}(a_{i})=\frac{1}{2}\left[\cos\left(\pi\times\frac{|a_{\text{pred}}-a_{\text{gt}}|-\epsilon_{1}\times|a_{\text{gt}}|}{(\epsilon_{2}-\epsilon_{1})\times|a_{\text{gt}}|}\right)+1\right],$ (6)
    where $a_{\text{pred}}$ is the predicted answer, $a_{\text{gt}}$ is the ground truth, $\epsilon_{1}$ is the tolerance threshold for an exact match (e.g., 0.05), and $\epsilon_{2}$ is the upper bound for partial rewards (e.g., 0.20). If $|a_{\text{pred}}-a_{\text{gt}}|<\epsilon_{1}\times|a_{\text{gt}}|$, the reward is 1 (exact match); if $|a_{\text{pred}}-a_{\text{gt}}|>\epsilon_{2}\times|a_{\text{gt}}|$, the reward is 0 (incorrect). This formulation ensures smooth transitions between full and partial rewards, enabling fair evaluation of numerical accuracy.
  • Function-based Type This reward type is designed for spatial transformation tasks requiring a sequence of transformation functions. The accuracy reward $R_{\text{acc}}(a_{i})$ evaluates the alignment between the predicted sequence $T_{\text{pred}}$ and the ground truth $T_{\text{gt}}$, computed as:
    $R_{\text{acc}}(a_{i})=\frac{\text{len}(T_{\text{pred}}^{f+o+v})+\alpha\cdot\text{len}(T_{\text{pred}}^{f+o/v})+\beta\cdot\text{len}(T_{\text{pred}}^{f})}{\max(\text{len}(T_{\text{pred}}),\text{len}(T_{\text{gt}}))},$ (7)
    where $T_{\text{pred}}^{f+o+v}$ is the subset of transformation steps with complete matches (w/ function, object, and value), $T_{\text{pred}}^{f+o/v}$ are the subsets with partial and only-function matches (w/ function and object, or w/ function and value), $T_{\text{pred}}^{f}$ is the subset with only-function matches. $\alpha$ and $\beta$ are the weighting coefficients for partial matches. This formulation ensures nuanced evaluation for flexible adjustment of partial match contributions.
4 Experiments
We design experiments to investigate the following key research questions:
  • RQ1: How effective is Reason-RFT in reasoning, generalization, and data efficiency?
  • RQ2: Why is the STAGE 1 of SFT with CoT reasoning necessary?
  • RQ3: Why is the STAGE 2 of reinforcement fine-tuning necessary?
  • RQ4: How does reward design affect Reason-RFT’s performance?
  • RQ5: What training dynamics emerge during reinforcement fine-tuning, and how do they shape the reasoning behavior of Reason-RFT?
4.1 Experimental Details
Datasets In this paper, we comprehensively evaluate the visual reasoning capabilities of our method by leveraging six existing datasets, enhanced through subtask categorization, error-prone data filtering, and dataset restructuring. Detailed protocols for data filtering and restructuring are provided in Sec. A. Specifically, we define three task categories as follows.
  • Visual Counting is a multimodal reasoning task evaluating the integration of linguistic, visual, and mathematical skills by solving arithmetic problems in 3D block-based scenes. Specifically, we filtered and corrected 35K samples from CLEVR-Math [1] for training and 1K test samples for in-domain (ID) evaluation. To assess generalization under domain-shift (DS), we constructed 1K new samples using 3D assets from Super-CLEVR [51], including two subsets: direct arithmetic (DS-D) and mixed arithmetic (DS-M). Refer to the Appendix Sec. A.1 for details.
  • Structure Perception is a structural reasoning task requiring models to analyze relationships in geometries, imaging structures, chart layouts, and architectural designs. We filtered 4.5K training samples and 820 ID test samples from Geo170K [52] and Math360K [55], along with 800 samples from Geometry3K [85] to evaluate DS adaptability. See the Appendix Sec. A.2.
  • Spatial Transformation is a spatial-visual reasoning task requiring models to infer single- or multi-step transformations by analyzing initial and final states of 3D scenes from multiple perspectives (e.g., center, left, right). We generated 100K samples using TRANCE [56], covering four difficulty levels, and selected 60K for training and 6K for testing through a specific filtering process. For DS evaluation, identical scenes are rendered from left/right viewpoints (DS-L/R) to test perspective-change robustness. Details can be found in the Appendix Sec. A.3.
Evaluation Metrics We use accuracy-rate (Acc) as the primary metric [86]. For numerical answers, correctness is verified by mathematical equivalence to the ground truth. For multiple-choice questions, we perform a string match. For function-type sequences, we use stepwise multi-level evaluation.
Implementation Details We utilize Qwen2-VL-2B and Qwen2-VL-7B [87] as the backbone models for our experiments. Our implementation is built on the open-source frameworks Open-R1 [88] and vLLM [89], ensuring reproducibility and scalability. All experiments were conducted on a cluster of servers, each equipped with $\timesA8$00 GPUs. For further details, see the Appendix Sec. B.
Training Paradigms and Baselines To assess the performance and generalization of different training strategies, we compare: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with CoT reasoning; and (2) RL-based methods—Reason-RFT-Zero, which applies RL without reasoning activation stage, and Reason-RFT, which uses limited CoT data for reasoning activation before RL training. For comprehensive experiments, we use Qwen2-VL-Instruct [87] as the base model (both 2B and 7B variants). In addition, we also select the most advanced open-source models  [90, 91, 92, 93, 94] and the proprietary models  [95, 96] as baselines to evaluate the performance of different paradigms.
4.2 Overall Evaluation of Reason-RFT Framework (RQ1)
To evaluate Reason-RFT, we evaluate Reason-RFT using 2B- and 7B-parameter models on three visual reasoning tasks. The results are summarized as follows.
Strong reasoning performance across ID tasks. As shown in Tab. 1, Reason-RFT achieves performance comparable to or better than both SFT- and RL-based methods across all tasks. In visual counting, Reason-RFT-Zero achieves the best performance among all models in the 7B setting. In structure perception, Reason-RFT outperforms most open-source and proprietary baselines in the 7B setting and remains competitive with top models such as InternVL-2.5-8B [92]. In spatial transformation, Reason-RFT matches or exceeds SFT-based methods while consistently outperforming all baselines. These results demonstrate that Reason-RFT effectively integrates the strengths of both SFT and RL in structured reasoning tasks.
Superior generalization under DS. Under DS settings, Reason-RFT shows substantial gains over both traditional baselines and alternative training paradigms. In visual counting, it outperforms ANS-SFT by 10.95% (2B) and 13.93% (7B). In structure perception, Reason-RFT achieves the highest performance in the 2B model, with an 6.93% gain over CoT-SFT, and remains highly competitive in the 7B model. Most notably, in the spatial transformation task, the 2B Reason-RFT model surpasses GPT-4o [95] by 31.7%, showcasing remarkable generalization under DS.
Figure 3: Results of different methods and model sizes on spatial transformation task across training.
High training efficiency. To evaluate data efficiency during training, we train all methods on the spatial transformation task and monitor intermediate ID and DS performance (Fig. 3). Additional results for visual counting and structure perception are provided in Appendix Sec. C. In the 2B model, Reason-RFT achieves 70% of the final performance of Reason-RFT-Zero using only 3% of the training data (1,600 samples), and reaches 82.5% with 9% of the data. In the 7B model, it achieves over 92% of Reason-RFT-Zero’s performance using just 3% of the data. These results confirm that Reason-RFT exhibits strong training efficiency in both ID and DS scenarios. The data-efficient nature of Reason-RFT renders it particularly effective for few-shot learning, offering significant potential for applications where labeled data is limited.
Table 1: Results on three visual reasoning tasks. The best results among different training paradigms are highlighted in bold, while the second-best results are underlined. “ID” denotes in-domain test data, and “DS” denotes domain-shift test data.
Method Visual Counting Structure Perception Spatial Transformation
ID DS-D DS-M AVG ID DS AVG ID DS-L DS-R AVG
Proprietary Models
GPT-4o-2024-08-06 [95] 68.10 42.54 9.60 40.08 50.18 43.49 46.83 42.55 28.67 29.76 35.88
Gemini-1.5-Pro [96] 61.80 41.20 26.40 43.13 50.12 48.38 49.45 26.22 18.76 19.88 22.77
Open-Source Models
Qwen2.5-VL-3B-Inst. [90] 75.90 50.93 4.40 43.74 36.75 37.44 37.09 8.57 8.26 8.31 8.42
Phi-3.5-Vision-4B-Inst. [91] 21.40 18.27 6.00 15.22 36.83 50.25 43.54 7.42 2.45 4.02 5.33
Qwen2.5-VL-7B-Inst. [90] 74.60 46.00 2.80 41.13 44.00 45.61 44.80 19.63 13.12 13.42 16.45
InternVL-2.5-8B [92] 93.50 46.13 2.80 47.48 63.00 47.32 51.60 7.19 6.62 6.63 6.91
Llama-3.2-11B-Vision [93] 10.30 9.60 9.20 9.70 13.75 20.85 17.30 8.22 8.40 9.03 8.47
Pixtral-12B [94] 42.60 25.33 15.60 27.84 30.38 36.09 33.23 7.35 5.03 5.22 6.42
Qwen2VL-2B-Instruct
Zero-Shot 82.40 42.67 0.00 41.69 25.86 20.63 23.25 3.78 4.60 4.67 4.35
++ ANS-SFT 96.20 51.07 5.20 50.82 51.34 22.50 36.92 77.39 49.24 50.33 58.99
++ CoT-SFT 85.50 49.73 36.80 57.34 43.05 25.25 34.15 64.37 43.19 42.86 50.14
++ Reason-RFT-Zero 98.40 58.00 5.20 53.87 47.68 32.50 40.09 42.13 34.07 33.41 33.74
++ Reason-RFT 96.80 60.00 28.40 61.77 49.03 33.13 41.08 74.61 64.05 64.08 67.58
Qwen2VL-7B-Instruct
Zero-Shot 98.60 54.53 4.80 52.64 43.30 43.88 43.59 13.53 12.72 12.78 13.01
++ ANS-SFT 95.00 42.53 8.00 48.51 51.34 25.38 38.36 82.19 54.29 54.83 63.77
++ CoT-SFT 87.30 45.33 33.60 55.41 50.49 33.00 41.75 81.31 47.90 47.80 59.00
++ Reason-RFT-Zero 99.40 63.60 21.20 61.40 55.00 54.75 54.88 67.67 57.20 56.15 60.34
++ Reason-RFT 95.60 56.13 35.60 62.44 59.27 49.25 54.26 79.97 59.36 58.61 65.98
4.3 Effect of STAGE 1 on Initialization (RQ2)
To investigate the role of CoT-SFT in initialization, we compare four baselines across three tasks. The results in Tab. 1 reveal the following: (1) Consistent performance gains from CoT-SFT. Across all three tasks and both 2B and 7B model scales, Reason-RFT consistently outperforms Reason-RFT-Zero following stage 1 reasoning activation. This improvement is particularly notable when the model is small and the task involves complex output structures. For example, in the spatial transformation task—which requires function-like serialized outputs—the 2B Reason-RFT model surpasses Reason-RFT-Zero by 33.84%. (2) Smaller models benefit more from CoT-SFT priors. In the visual counting task under the DS-M setting, the 2B model with CoT-SFT outperforms Reason-RFT-Zero by 31.6%. Although the gap narrows in the 7B model, CoT-SFT still yields substantial gains. This indicates that pure RL-based methods struggle to adapt from direct arithmetic to mixed arithmetic reasoning under DS, whereas CoT-SFT provides effective inductive priors for such adaptation. Moreover, under the same amount of CoT-SFT data, the 2B model still underperforms its 7B counterpart, highlighting the increased reliance of smaller models on CoT-SFT for acquiring reasoning capabilities.
Figure 4: Results of DS v.s. ID on spatial transformation task.
4.4 Effect of STAGE 2 on Generalization (RQ3)
To evaluate the impact of RL in stage 2, we compare the generalization performance of CoT-SFT and Reason-RFT across three visual reasoning tasks under DS. As shown in Tab. 1 and Fig. 4, the results reveal the following: Reinforcement fine-tuning improves generalization beyond CoT-SFT. Across all domain-shift settings, Reason-RFT consistently outperforms CoT-SFT, demonstrating that reinforcement learning significantly enhances model robustness. For instance, in the visual counting task, the 7B Reason-RFT achieves a combined DS-D and DS-M score 12.8% higher than CoT-SFT. The improvement is even more pronounced in structure-sensitive tasks such as spatial transformation, where the 2B Reason-RFT exceeds CoT-SFT by 21.04% on average across DS-L and DS-R. These results indicate that CoT-SFT alone yields limited generalization, while reinforcement fine-tuning enables better adaptation to compositional and layout-dependent variations.
4.5 Exploration on Reward Design (RQ4)
Table 2: Results of different format reward strategies on the spatial transformation task.
Setting ID DS-L DS-R AVG
Qwen2VL-2B-Instruct
Reason-RFT-Zero 42.13 34.07 33.41 33.74
++ visual tokens 42.01 36.05 35.97 38.01
Reason-RFT 74.61 64.05 64.08 69.33
++ visual tokens 71.99 60.13 59.87 65.99
Qwen2VL-7B-Instruct
Reason-RFT-Zero 67.67 57.2 56.15 62.17
++ visual tokens 70.28 59.52 57.01 64.27
Reason-RFT 79.97 59.36 58.61 69.48
++ visual tokens 79.85 58.71 57.98 69.09
Format Reward. In DeepSeek-R1 [39], the format reward enforces the use of <think> and <answer> tokens to structure reasoning in textual tasks. To better support visual reasoning, we extend this with <summary> and <caption> tokens to incorporate visual observations via caption-style prompts. As shown in Tab. 2, this improves Reason-RFT-Zero but has limited effect on Reason-RFT. We attribute this to Reason-RFT’s prior CoT supervision, which likely helps it internalize caption-like structures in stage 1, reducing the benefit of explicit tags. In contrast, Reason-RFT-Zero benefits more from such structural cues, indicating greater sensitivity to format-level guidance.
Table 3: Results of different accuracy reward strategies on the spatial transformation task.
Setting $\alpha$ $\beta$ ID DS-L DS-R AVG
Qwen2VL-2B-Instruct
Baseline 0 0 74.61 64.05 64.08 69.33
(a) 0.50 0.25 79.18 56.36 55.45 67.54
(b) -0.25 -0.50 73.69 64.41 64.72 69.13
Qwen2VL-7B-Instruct
Baseline 0 0 79.97 59.36 58.61 69.48
(a) 0.50 0.25 80.89 53.20 52.61 66.90
(b) -0.25 -0.50 75.03 64.83 63.18 69.52
Accuracy Reward. We explore accuracy reward design in the spatial transformation task, which requires predicting transformation sequences in a structured format. The formulation in Eq. 7 introduces coefficients $\alpha$ and $\beta$ to control tolerance for partial matches. We test three settings: (1) $\alpha=0,\beta=0$ (exact match only), (2) $\alpha=0.50,\beta=0.25$ (partial credit), and (3) $\alpha=-0.25,\beta=-0.50$ (penalized partial matches). Results on 2B and 7B models (Tab. 3) show that: (1) partial credit improves ID performance but harms generalization, suggesting “soft rewards” reduce robustness; (2) penalizing partial matches improves generalization under domain shift, indicating “hard rewards” better support serialized reasoning.
Figure 5: Analysis of Greedy Reward Stratification. The model’s reasoning token length first decreases, then gradually rises and stabilizes during Reason-RFT-Zero training. The peak of the format reward coincides with the accuracy reward’s rapid growth phase.
Figure 6: Illustration of the Transient Adaptation Gap. (a) shows a sharp drop and recovery in both ID and DS test performances within the early training steps. (b) shows a case study of the prediction result on early step.
Figure 7: Analysis of Reasoning Redundancy. (a) shows the reasoning token length curves for Reason-RFT-Zero and Reason-RFT during training. (b) displays their accuracy reward, with both paradigms converging to similar accuracy rate.
4.6 Training Dynamics and Reasoning Behavior Analysis (RQ5)
Greedy Reward Stratification. This phenomenon captures the model’s tendency, particularly in Reason-RFT-Zero, to prioritize easier-to-optimize rewards (e.g., format reward) before addressing more challenging objectives (e.g., accuracy reward). As shown in Fig. 5, the reasoning token length initially drops, then gradually increases and stabilizes. This dynamic correlates with the format reward quickly reaching a plateau, followed by a sharp rise in the accuracy reward. We infer that the model initially simplifies its outputs to adapt rapidly to structured format expectations, and only later allocates learning capacity to improve reasoning correctness.
Transient Adaptation Gap. This refers to a temporary performance degradation observed in the early training phase of Reason-RFT-Zero. When the model shifts from directly predicting answers to producing structured reasoning traces, it experiences a brief adaptation bottleneck—marked by a sharp decline and gradual recovery in accuracy. Fig. 7 (a) illustrates this drop within the first 100 steps on the visual counting task under both ID and DS settings. A case analysis in Fig. 7 (b) further reveals that forcing structured reasoning prematurely may lead to incorrect outputs, highlighting the sensitivity of early-stage training to reasoning format constraints.
Reasoning Redundancy. This phenomenon concerns the discrepancy in reasoning token length between models trained with and without CoT activation. In the structure perception task, Reason-RFT and Reason-RFT-Zero attain similar accuracy, yet the former generates significantly longer reasoning traces (Fig. 7). This likely stems from Reason-RFT’s use of CoT data distilled from advanced models (e.g., GPT-4o), which encourages verbose reasoning during stage 1. In the absence of penalties or length control in reinforcement fine-tuning, such verbosity persists. By contrast, Reason-RFT-Zero converges to more concise reasoning through reward-driven exploration. We hypothesize that these longer chains in Reason-RFT may introduce unnecessary computational overhead or reflect overthinking relative to task complexity.
5 Conclusion
In this paper, we propose Reason-RFT, a novel reinforcement fine-tuning framework that enhances the generalization capabilities of visual reasoning models. By integrating SFT with CoT reasoning activation data and GRPO-based reinforcement learning, Reason-RFT effectively mitigates key challenges such as overfitting and cognitive rigidity, thereby improving cross-domain transferability and real-world applicability. To support systematic evaluation, we reconstruct a comprehensive dataset covering visual counting, structure perception, and spatial transformation tasks, establishing a robust benchmark for assessing model performance across diverse scenarios. Extensive experiments demonstrate the effectiveness of Reason-RFT, providing valuable insights for advancing visual reasoning research and introducing a new paradigm in multimodal learning.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62476011), and the National Science and Technology Major Project (No. 2022ZD0117800).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Appendix Section H Limitations and Societal Impact.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results or proofs but focuses on architectural innovation and empirical evaluation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Section 4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: See supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Section 4 Implementation details and Appendix Sec. B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We use five various seeds to train the model and report the average accuracy of them to avoid randomness.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Section 4 Implementation details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read the NeurIPS Code of Ethics, and our paper does not have these problems.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Appendix H Limitations and Societal Impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: CC-BY 4.0.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: The core methodology and experiments do not involve LLMs. Any language editing assistance does not impact the scientific contributions.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
This supplementary material provides additional details on the proposed method and experimental results that could not be included in the main manuscript due to page limitations. Specifically, this appendix is organized as follows.
  • Sec. A provides more details on the evaluation of reasoning tasks and discusses how we collected, filtered, and reconstructed a high-quality dataset.
  • Sec. B outlines the models and training processes, providing more detailed experimental specifics.
  • Sec. C presents comprehensive experimental results.
  • Sec. D details the pipeline of CoT date generation.
  • Sec. E presents detailed composition of diffenernt mixed CoT datasets.
  • Sec. F shows the comparison of CoT quality before and after RL.
  • Sec. G includes more visualization cases.
  • Sec. H introduces the limitations of our Reason-RFT and its societal impact.
Appendix A Details of Evaluation Reasoning Tasks
A.1 Visual Counting
Figure 8: The sample of Visual Counting.
Task Definition Visual Counting is a multi-modal reasoning task that evaluates the integration of linguistic, visual, and mathematical capabilities by requiring models to solve arithmetic problems in dynamic visual scenes composed of 3D blocks with diverse attributes, including color, size, material, and shape. The task consists of four distinct reasoning types: 1) Subtraction, which involves counting objects after removing a specified subset based on given attributes; 2) Addition, where models must compute totals after inserting new objects with defined quantities and properties; 3) Adversarial, a challenging variant designed as trick questions in which operations are performed on one set of objects while the query targets an unrelated or unaffected subset, testing the model’s robustness against deceptive scenarios; and 4) Multi-Hop, which requires sequential reasoning through multiple addition or subtraction steps to arrive at the final count. This task challenges models to perform attribute-based reasoning in dynamic visual contexts, emphasizing cross-modal understanding and reasoning capabilities. Some examples are shown in Fig. 8.
Dataset Preparation For In-Domain (ID) dataset, we refined the original dataset from CLEVR-Math [1] by filtering out low-quality or incorrect samples using GPT-4o, resulting in a clean dataset comprising 35K training samples and 1K test samples. These samples are categorized into four specific types: subtraction, addition, adversarial, and multihop-subtraction. To evaluate Domain-Shift (DS) generalization, we extended CLEVR-Math by enhancing the diversity of objects through the incorporation of 3D assets from Super-CLEVR [51], which leads to the creation of Super-CLEVR-Math, an advanced benchmark with 1K test samples designed to assess model generalization under increased complexity. These test samples are also divided into four task types: addition, subtraction, subtraction-multihop and addition-subtraction. The test samples are further categorized into four task types: addition, subtraction, subtraction-multihop, and addition-subtraction. The first three constitute the DS-D subset, while addition-subtraction forms the DS-M subset. Notably, the mixed addition-subtraction type introduces a novel category consisting multi-steps of both addition and subtraction, which is not present in CLEVR-Math, further elevating the benchmark’s challenge.
Reward Design Following the reward methodology of DeepSeek-R1 [39], we define two distinct reward functions: Format Reward and Accuracy Reward. The Format Reward is assigned a value of 1 if the response adheres to the predefined template structure, specifically in the form of <think>...</think><answer>...</answer>; otherwise, it is assigned a value of 0. The Accuracy Reward is assigned a value of 1 if the numerical counting result in the response is correct; otherwise, it is assigned a value of 0. This dual-reward mechanism ensures both structural compliance and numerical accuracy in model responses.
A.2 Structure Perception
Figure 9: The sample of Structure Perception.
Task Definition Structure Perception represents a complex class of visual mathematical reasoning tasks, which focuses on assessing the model’s capacity to determine geometric structure relationships and perform calculations involving angles, lengths, areas, and other geometric properties. The task includes problems such as identifying congruent or similar shapes, calculating perimeters and areas, determining angles between lines or shapes, and solving problems related to geometric transformations (e.g., rotations, translations, and reflections). By combining mathematical rigor with visual reasoning, this task challenges models to demonstrate a deep understanding of geometric principles in both abstract and real-world scenarios. Some examples are shown in Fig. 9.
Dataset Preparation For the ID dataset, we provide GeoMath-4K5, a dataset specifically designed for geometric problem solving, which is constructed based on Math360K [55] and Geo170K [52]. To ensure data quality, we employed GPT-4o to filter out incorrect samples and removed those with answers that were neither numerical nor included in the provided options, thereby streamlining the validation process during training and testing. This refinement process resulted in a curated dataset consisting of 4.5K training samples and 820 test samples. For DS evaluation, we selected 800 samples from Geometry3K [85] (including 400 multiple-choice and 400 open-ended questions) to comprehensively assess the model’s generalization capabilities on geometry reasoning.
Reward Design We maintain the same Format Reward as used in the Visual Counting task above. The Accuracy Reward is extended to support the evaluation of both multiple-choice questions and mathematical expressions, ensuring comprehensive assessment across various problem types. Specifically, mathematical reward type is designed for Structure Perception tasks involving numerical answers, such as floating-point values or LaTeX-formatted expressions. It uses a tolerance-based evaluation to account for minor numerical deviations. The accuracy reward $R_{\text{acc}}(a_{i})$ is defined as:
$R_{\text{acc}}(a_{i})=\frac{1}{2}\left[\cos\left(\pi\times\frac{|a_{\text{pred}}-a_{\text{gt}}|-\epsilon_{1}\times|a_{\text{gt}}|}{(\epsilon_{2}-\epsilon_{1})\times|a_{\text{gt}}|}\right)+1\right],$ (8)
where $a_{\text{pred}}$ is the predicted answer, $a_{\text{gt}}$ is the ground truth, $\epsilon_{1}$ is the tolerance threshold for an exact match (e.g., 0.05), and $\epsilon_{2}$ is the upper bound for partial rewards (e.g., 0.20). If $|a_{\text{pred}}-a_{\text{gt}}|<\epsilon_{1}\times|a_{\text{gt}}|$, the reward is 1 (exact match); if $|a_{\text{pred}}-a_{\text{gt}}|>\epsilon_{2}\times|a_{\text{gt}}|$, the reward is 0 (incorrect). This formulation ensures smooth transitions between full and partial rewards, enabling fair numerical evaluation.
A.3 Spatial Transformation
Task Definition Spatial Transformation is a spatial-visual reasoning task designed to infer single-step or multi-step transformation actions by analyzing the initial and final visual states from multiple perspectives (e.g., center, left, right). The task utilizes transformation functions, including change_size, change_color, change_material, change_shape, and change_position, to modify object properties such as size, color, material, shape, and position using predefined values. This task evaluates the model’s ability to reason about spatial relationships and object transformations across diverse viewpoints in dynamic visual scenarios. Some examples are shown in Fig. 10.
Dataset Preparation We generated 100K samples using the environment and configuration from Trance [56], with each sample comprising initial object attributes, front-view image of initial state, and images of final state captured from front, left, and right perspectives. To ensure high data quality, we implemented a rigorous filtering process: (1) removing samples containing occluded or invisible objects in either the initial or final states, (2) eliminating redundant actions within the transformation sequences, and (3) consolidating multi-step displacement actions, which collectively ensure the uniqueness and correctness of the solutions. The refined dataset consists of 60K training samples and 6K test samples. For the training set, we constructed the Trans-Center-60K dataset using the Center-Center configuration, which pairs front-view initial and final state images. For ID evaluation, we derived the Trans-Center-6K dataset from the 6K test samples under the same Center-Center configuration. To evaluate DS generalization, we constructed two additional datasets: Trans-Left-6K (DS-L) and Trans-Right-6K (DS-R), leveraging the Center-Left and Center-Right configurations to assess the model’s generalization capabilities in spatial reasoning under viewpoint conditions.
Figure 10: The sample of Spatial Transformation.
Reward Design For the Format Reward, we adopted the same formulation as used in the Visual Counting task. As for the Accuracy Reward, a specialized design was developed to evaluate the sequence of transformation functions. Function-based type is designed for Spatial Transformation tasks requiring a sequence of transformation functions. The accuracy reward $R_{\text{acc}}(a_{i})$ evaluates the alignment between the predicted sequence $T_{\text{pred}}$ and the ground truth $T_{\text{gt}}$, computed as:
$R_{\text{acc}}(a_{i})=\frac{\text{len}(T_{\text{pred}}^{f+o+v})+\alpha\cdot\text{len}(T_{\text{pred}}^{f+o/v})+\beta\cdot\text{len}(T_{\text{pred}}^{f})}{\max(\text{len}(T_{\text{pred}}),\text{len}(T_{\text{gt}}))},$ (9)
where $T_{\text{pred}}^{f+o+v}$ is the subset of transformation steps with complete matches (w/ function, object, and value), $T_{\text{pred}}^{f+o/v}$ are the subsets with partial and only-function matches (w/ function and object, or w/ function and value), $T_{\text{pred}}^{f}$ is the subset with only-function matches. $\alpha$ and $\beta$ are the weighting coefficients for partial matches. This formulation ensures nuanced evaluation, rewarding both exact and partially correct responses while allowing flexible adjustment of partial match contributions.
System Prompts For the Spatial Transformation task, we designed two versions of the system prompt. The first version specifies the answer output format using the <think> and <answer> tags, while the second version includes additional outputs <summary> and <caption> for experiments on exploration of format reward design in the main paper. These two versions are illustrated in Fig. 11 and Fig. 12, respectively.
Appendix B Details of Models and Training
We utilize Qwen2-VL-2B and Qwen2-VL-7B [87] as the backbone models for our experiments. Our implementation is built on the open-source frameworks Open-R1 [88] and vLLM [89], ensuring reproducibility and scalability. All experiments were conducted on a cluster of servers, each equipped with $\timesA8$00 GPUs. For the Visual Counting task and Spatial Transformation task, we trained the models for 1 epoch each on their respective training datasets, ensuring sufficient exposure to task-specific patterns while avoiding overfitting. For the Structure Perception task, due to its GeoMath training dataset consisting of a relatively small number of training samples (a total of 4,500), we extended the training duration to 5 epochs to allow the models to fully capture the underlying structural and geometric relationships. In the Reason-RFT training pipeline, all models underwent an initial CoT activation stage with 1,600 samples before proceeding to the RL phase. More details on training process of each models are shown in Tab. 4
Table 4: Detailed configuration for each training stage of Reason-RFT. The table presents the training parameters for the 2B model and 7B model across three visual reasoning tasks. The parameters marked with correspond to Visual Counting / Structure Perception / Spatial Transformation.
Qwen2-VL-2B Qwen2-VL-7B
Stage-1 Stage-2 Stage-1 Stage-2
Data
Dataset CoT dataset Non-CoT dataset CoT dataset Non-CoT dataset
#Samples 1.6K 35K / 4.5K / 60K 1.6K 35K / 4.5K / 60K
Model
Trainable Part Full Model Full Model Full Model Full Model
#Tunable Parameters 2.21B 2.21B 8.29B 8.29B
Training
Per-device Batch Size 8 1 8 1
Gradient Accumulation 2 2 2 2
LR: $\{\psi_{v}^{\text{ViT}},\phi_{v}^{\text{LLM}}\}$ $\times 10^{-5} 1$ $\times 10^{-6}$ 1 $\times 10^{-5}$ 1 $\times 10^{-6}$
Epoch 1 1 / 5 / 1 1 1 / 5 / 1
Optimizer AdamW AdamW AdamW AdamW
Deepspeed Zero3 Zero3 Zero3 Zero3
Weight Decay 0.1 0.0 0.1 0.0
Warmup Ratio 0.03 0.00 0.03 0.00
LR Schedule Cosine Cosine Cosine Cosine
Max Seq. Length 32768 4096 32768 4096
Max Compl. Length 512 512
Num. of Compl. 8 4
GPU Nums 1 $\times$ 8 1 $\times$ 8 1 $\times$ 8 1 $\times$ 8
Appendix C More Experiment Results
Table 5: Results of various mixed CoT activation datasets on the Structure Perception task.
Setting CoT Activation Data ID DS AVG
Baseline GeoMath-only data 59.27 49.25 54.26
(a) Mixed Specific-Domain data 50.61 45.35 48.02
(b) Mixed General-Domain data 42.51 40.25 41.38
Exploration on COT Activation Data To investigate the impact of differently composed CoT activation data on Reason-RFT, we construct two distinct datasets: a mixed domain-specific dataset, which integrates relevant yet distinct data from in-domain tasks, and a mixed general-domain dataset, encompassing a broader range of visual reasoning tasks (e.g., graph topology, visual puzzles). The detailed dataset composition is provided in Appendix Sec. E. Using these datasets, we perform Reason-RFT training on Structure Perception task, with the results detailed in Tab. 5. From this, two key points emerge: (1) As the proportion of in-domain training data decreases, the model’s performance on specific tasks declines; (2) Models trained on more diverse visual reasoning domain data may also exhibit a reduction in domain-specific performance.
Results on Different Backbones We further validate the effectiveness of Reason-RFT on stronger or alternative vision–language backbones. We report results on three visual reasoning tasks: Visual Counting (T1), Structure Perception (T2), and Spatial Transformation (T3) in the combined Tab. 6. Reason-RFT achieves the strongest averages across backbones and tasks, with especially large margins on domain-shifted splits while keeping in-domain (ID) performance near ceiling. On Qwen2.5-VL-3B, for T1 (Visual Counting) ID is already saturated (99.0 with Reason-RFT-Zero vs. 98.8 with Reason-RFT), yet Reason-RFT markedly improves robustness on DS: +9.2+\!9.2 on DS-D (68.7 vs. 59.5 vs. CoT-SFT) and +5.6+\!5.6 on DS-M (54.8 vs. 49.2 vs. CoT-SFT), with an especially large +44.0+\!44.0 over Reason-RFT-Zero on DS-M (54.8 vs. 10.8), yielding the best T1 AVG (74.1); for T2 (Structure Perception) it is best on both ID/DS (59.0/56.6), beating CoT-SFT by +2.9+\!2.9 (ID) and +7.2+\!7.2 (DS) and Reason-RFT-Zero by +4.2+\!4.2 (ID) and +2.1+\!2.1 (DS), indicating that RL enhances stepwise structural reasoning rather than overfitting; for T3 (Spatial Transformation) ANS-SFT attains the top ID (91.1) but is less robust, whereas Reason-RFT trades a modest 4.4-4.4 on ID (86.7) for substantial DS gains of +8.2+\!8.2 on DS-L (55.2 vs. 47.0) and +7.6+\!7.6 on DS-R (54.4 vs. 46.8), delivering the best AVG (65.4) and a superior ID/DS Pareto. On InternVL3-2B for T1, Reason-RFT is best on ID/DS-D/DS-M (99.10/69.80/55.90) and AVG (74.93), improving over ANS-SFT, CoT-SFT, and Reason-RFT-Zero by +23.16+\!23.16, +8.53+\!8.53, and +18.06+\!18.06, respectively, with the largest domain-shift margin on DS-M of +43.60+\!43.60 over Reason-RFT-Zero (55.90 vs. 12.30). Taken together, these trends across architecturally distinct backbones indicate that the benefits of Reason-RFT are backbone-agnostic, improving both coherence-driven reasoning and out-of-distribution reliability.
Table 6: Results on different backbones across three tasks. Best is bold; second-best is underlined. “ID” denotes in-domain; “DS-*” denotes domain-shifted splits. Missing results are shown as “–”.
Backbone Method Visual Counting (T1) Structure Perception (T2) Spatial Transformation (T3)
ID DS-D DS-M AVG ID DS AVG ID DS-L DS-R AVG
Qwen2.5-VL-3B Zero-Shot 75.9 50.9 4.4 43.7 36.8 37.4 37.1 8.6 8.3 8.3 8.4
++ ANS-SFT 97.4 51.5 6.0 51.6 53.0 31.8 42.4 91.1 47.0 46.8 61.6
++ CoT-SFT 89.2 59.5 49.2 66.0 56.1 49.4 52.7 81.6 46.1 44.2 57.3
++ Reason-RFT-Zero 99.0 58.9 10.8 56.2 54.8 54.5 54.6 68.5 49.5 48.0 55.3
++ Reason-RFT 98.8 68.7 54.8 74.1 59.0 56.6 57.8 86.7 55.2 54.4 65.4
InternVL3-2B Zero-Shot 79.30 51.20 5.10 45.20
++ ANS-SFT 96.80 52.00 6.50 51.77
++ CoT-SFT 88.90 60.10 50.20 66.40
++ Reason-RFT-Zero 98.90 59.40 12.30 56.87
++ Reason-RFT 99.10 69.80 55.90 74.93
Table 7: Evaluation results on general benchmarks.
Method General
MMMU RealWorldQA MathVision AI2D
Qwen2VL-2B-Instruct
Zero-Shot 39.89 61.31 12.50 72.50
++ ANS-SFT 40.56 48.76 15.79 68.20
++ CoT-SFT 34.00 37.78 12.99 65.36
++ Reason-RFT-Zero 39.30 42.81 13.00 74.61
++ Reason-RFT 41.14 53.06 14.82 75.24
Qwen2VL-7B-Instruct
Zero-Shot 54.10 67.19 16.30 83.00
++ ANS-SFT 42.66 48.10 9.12 78.30
++ CoT-SFT 44.67 36.46 15.30 73.25
++ Reason-RFT-Zero 46.44 45.10 10.86 75.28
++ Reason-RFT 50.04 61.31 17.60 81.70
Evaluation on General Benchmarks Although Reason-RFT is primarily designed to enhance domain-specific visual reasoning abilities, we conduct a thorough evaluation on general benchmarks to verify whether our approach compromises the model’s general reasoning capabilities. Tab. 7 presents the results on four widely adopted datasets: MMMU [97], RealWorldQA [98], MathVision [99], and AI2D [100]. Across all tasks and model scales, Reason-RFT consistently maintains or even improves general performance. For instance, on the 2B model, Reason-RFT achieves the highest scores on MMMU (41.14) and AI2D (75.24), outperforming both zero-shot baselines and other supervised fine-tuning approaches such as ANS-SFT and CoT-SFT. Notably, it also improves performance on the challenging MathVision task (14.82), demonstrating its robustness in spatial reasoning. For the larger 7B model, Reason-RFT again surpasses ANS-SFT and CoT-SFT by large margins, particularly on RealWorldQA (61.31) and MMMU (50.04), while maintaining strong results on AI2D (81.70). These results suggest that Reason-RFT not only scales effectively with model size but also introduces no observable performance degradation on general benchmarks. In summary, the empirical evidence supports that Reason-RFT enhances domain-specific reasoning while preserving—if not enhancing—general visual-language reasoning capabilities. This confirms the robustness and transferability of our method, making it a strong alternative to conventional fine-tuning paradigms.
Performance at Different Training Steps Fig. 13 and Fig. 14 illustrate the ID and DS performance of all the training methods across three visual reasoning tasks, evaluated at various training sample sizes. This analysis helps us understand how each method scales with training data. More detail evaluation results for each subset of three tasks are in Tab. 10 - Tab. 17. We systematically varied the number of training samples, from minimal to substantial, allowing us to identify performance thresholds and data efficiency for each method in both ID and DS contexts. Key findings from this analysis include: Data Efficiency of Reason-RFT: Reason-RFT demonstrates exceptional data efficiency, achieving approximately 70% of the performance of Reason-RFT-Zero with only 3% of the training data (1,600 samples), and 82.5% with just 9%. Robust Generalization to DS scenarios: In the 7B model, Reason-RFT achieves over 92% of Reason-RFT-Zero’s performance using just 3% of the training data, showcasing its strong generalization capabilities. Comparison Across Methods: Reason-RFT consistently outperforms other methods, particularly in data-constrained scenarios, indicating its suitability for applications with limited data availability. Performance Saturation: As training sample size increases, some methods experience performance plateaus, suggesting that beyond a certain point, additional data yields diminishing returns.
In conclusion, the evaluation of performance across different training samples not only highlights the strengths of Reason-RFT in terms of data efficiency and generalization but also provides critical insights into the performance dynamics of various methods. These findings are essential for practitioners aiming to maximize performance while effectively managing training resources.
Appendix D More Details on CoT Data Construction
This section expands the pipeline of CoT generation by detailing both the automated and manual components used to construct our CoT-SFT corpus.
(1) Automated Generation. We instantiate CoT drafts using reasoning-guided prompt templates such as “Let’s break down the problem step by step…” and “To answer this, we need to consider…”. Templates are combined with model prompting (GPT-4o [95] and Gemini-Pro [96]) under temperature-controlled sampling $(T{=}0.7,~\text{top-}k{=}50,~\text{top-}p{=}0.9)$. To increase coverage and depth, we insert hand-crafted, subtask-specific few-shot exemplars that bias toward explicit intermediate justifications and error-checking behavior.
(2) Automated Filtering. Each generated CoT is screened by two criteria:
Length range. For each subtask $s$, we compute a target trajectory length $\bar{L}_{s}$ from a balanced mixture of 50% human-written and 50% model-generated samples:
$\bar{L}_{s}\;=\;\tfrac{1}{2}\big(\bar{L}^{\text{human}}_{s}+\bar{L}^{\text{model}}_{s}\big),$
where $\bar{L}^{\text{human}}_{s}$ and $\bar{L}^{\text{model}}_{s}$ are computed as sample means over their respective sets after basic de-duplication. A candidate with length $L_{i}$ is retained iff
$0.6\,\bar{L}_{s}\;\leq\;L_{i}\;\leq\;1.4\,\bar{L}_{s},$
where $L_{i}$ is measured in tokens by our training-time tokenizer; the factors $0.6$ and $1.4$ were selected via a small pilot study to trim outliers while preserving diversity.
Inconsistency. We discard the responses that contradict the known ground truth, including self-inconsistent counts, incompatible algebraic steps, or reasoning that invalidates later conclusions.
For reference, the empirical trajectory-length statistics (mean $\mu$, stdev $\sigma$) across tasks are:
Task $\mu$ (tokens) $\sigma$ (tokens)
Visual Counting 70 30
Structural Perception 180 80
Spatial Transformation 400 120
where $\mu$ and $\sigma$ are computed over the curated pool after automated filtering and before manual review.
(3) Human Verification. We randomly sample $10\%$ of CoT drafts from eachtask for manual review, focusing on (i) step-to-step coherence, (ii) logical validity, and (iii) alignment between the reasoning chain and the final answer. Typical failure modes include: (i) a correct final answer supported by an incorrect chain (e.g., deriving triangle area via the Pythagorean theorem); (ii) internal contradictions, such as stating “There are 3 red blocks on the left and 2 on the right” and later concluding the total is 6. A follow-up quality audit found that, prior to human verification, approximately $3.8\%$ of samples contained critical logical flaws; after verification, the residual error rate fell below $1\%$, indicating high post-cleanup reliability.
Discussion. Automated generation with calibrated sampling and subtask-specific few-shots provides diverse yet structured CoTs; the length- and consistency-based filters remove overly terse/verbose or self-contradictory drafts; targeted human verification further suppresses high-severity errors. Together, these stages yield a CoT–SFT dataset with improved coherence and faithfulness, while maintaining scalability and reproducibility.
Appendix E Detail on Mixed CoT Datasets
As shown in Tab. 8, we presents a comprehensive overview of the datasets utilized for all of our visual reasoning experiments, categorized into three experimental groups. All of them are CoT-annotated by GPT-4o [95] The Main Experiment section includes three large-scale datasets: Visual-Counting (35,000 samples) for quantitative analysis, Structure-Perception (4,500 samples) for structural understanding, and Spatial-Transformation (60,000 samples) assessing spatial reasoning capabilities. For Ablation Studies, two mixed-domain subsets were constructed: (1) The Mixed General-Domain set comprises 11 CoT-annotated datasets spanning scientific reasoning (AI2D [59], ScienceQA [101]), topological graph problems (GVLQA series [102]), and pattern recognition (PuzzleVQA [103], IconQA [104], Raven [105]). (2) The Mixed Specific-Domain set focuses exclusively on geometric reasoning, featuring GeoQA [106], GeomVerse [53], and Geometry3K [85] with progressively complex problem structures. All datasets were standardized to ensure training compatibility.
Table 8: Datasets Overview for Visual Reasoning Tasks
Dataset Name Samples Reasoning Type Description
Main Experiment
Visual-Counting 35,000 Visual Counting Full dataset for visual counting task
Structure-Perception 4,500 Structure Perception Full dataset for structural perception tasks
Spatial-Transformation 60,000 Spatial Transformation Full dataset for spatial transformation tasks
Ablation Experiment (Mixed General-Domain)
AI2D [59] 1,467 Scientific Reasoning Scientific diagram interpretation
ScienceQA [101] 2,112 Scientific Reasoning Science question answering
GVLQA-connectivity [102] 1,199 Topological Reasoning Graph connectivity problems
GVLQA-cycle [102] 1,194 Topological Reasoning Cycle detection in graphs
GVLQA-hamilton [102] 1,158 Topological Reasoning Hamiltonian path problems
GVLQA-topology [102] 1,070 Topological Reasoning General topology questions
GVLQA-matching [102] 1,193 Topological Reasoning Graph matching tasks
PuzzleVQA [103] 1,618 Pattern/Puzzle Visual puzzle solving
IconQA [104] 5,270 Pattern/Puzzle Icon-based question answering
Raven [105] 982 Pattern/Puzzle Raven’s Progressive Matrices
Ablation Experiment (Mixed Specific-Domain)
GeoQA [106] 1,500 Geometric Reasoning Geometric problem solving
GeomVerse [53] 2,841 Geometric Reasoning Advanced geometry challenges
Geometry3K [85] 3,794 Geometric Reasoning Comprehensive geometry problems
Appendix F Comparison of CoT Quality Before and After RL
Setting. We compare the Qwen2VL-3B model trained with Reason-RFT (Stage 2, post-RL) against the same backbone trained with only CoT-SFT (Stage 1, pre-RL) on the Structure Perception task. Unless otherwise noted, statistics are computed over a random sample of $n=100$ problem instances.
Table 9: Summary of comparative metrics on Structure Perception ($n=100$). Positive values indicate post-RL improvements.
Metric Change (Post–Pre)
Reasoning Step Count +2.7 steps
Prompting Words +14%
Logical Connectives +23%
Answer Accuracy +20.56%
Qualitative findings. Despite the high textual similarity between the two variants, the post-RL model exhibits stronger logical coherence across intermediate steps, with fewer broken or skipped chains of inference. For example, in Fig. 16 (case 2), the pre-RL model correctly infers a formula but omits the subsequent multiplication by $2$, an error that is notably less frequent after RL. In addition, the post-RL model more often displays reflective behaviors (e.g., “let me double check”) that are rarely observed in pre-RL outputs as shown in the math example of Fig. 1.
Quantitative protocol. We assess three dimensions of chain-of-thought (CoT) quality: (i) Reasoning Step Count—the number of explicitly delimited reasoning steps per sample, obtained via automatic counting with GPT-4o [95]; (ii) Lexical Usage of two categories of expressions: Prompting words (e.g., “oh I see”, “let me think step by step”, “let me double check”) and Logical connectives (e.g., “so”, “therefore”, “first”, “but”, “moreover”); and (iii) Answer Accuracy as reported for the accuracy rate of Structure Perception task.
Results and interpretation. Post-RL training increases the average CoT granularity (as reflected by the larger step count), the organizational scaffolding of reasoning (higher usage of prompting phrases and discourse connectives), and the task effectiveness (higher final-answer accuracy). Taken together, these observations indicate that reinforcement learning with Reason-RFT enhances both the coherence and utility of CoT: it reduces fragile or truncated chains, encourages reflective self-checks, and translates these behaviors into substantial accuracy gains.
Appendix G Visualization
In this section, we present additional visualization results on general visual reasoning and three specific task reasoning, see Fig. 15 - Fig. 23. Reason-RFT demonstrates superior performance over CoT-SFT in terms of logical consistency, reasoning quality, and correctness. CoT-SFT’s flaws stem from incorrect assumptions and misinterpretations, highlighting the importance of accurate problem interpretation and reasoning in visual reasoning tasks.
Appendix H Limitations and Societal Impact
Limitations While Reason-RFT has demonstrated strong performance in visual reasoning tasks, there are still areas to address. Future work will explore its application across a range of computer vision models, scaling to larger architectures (e.g., 32B/72B), and integrating large-scale Mixture of Experts (MoE) models to evaluate generalization. We will also extend the framework to complex downstream scenarios, such as embodied AI and autonomous driving, testing its effectiveness in real-world applications that require sophisticated visual reasoning and real-time decision-making.
Societal Impact The advancements of Reason-RFT in visual reasoning have important societal implications. By enhancing generalization and cross-domain transferability, this framework can improve AI applications in areas like medical imaging, autonomous driving, and assistive technologies for the visually impaired. It also reduces overfitting and cognitive rigidity, leading to more reliable and interpretable AI systems that foster trust in human-AI collaboration. The reconstructed benchmark dataset allows for fair evaluation, promoting research in robust AI. However, ethical considerations, such as biases in training data and responsible deployment, must be addressed to prevent misuse. Overall, Reason-RFT paves the way for adaptable and trustworthy AI, benefiting industries, researchers, and society.
Figure 11: The system prompt used in Spatial Transformation task.
Figure 12: The system prompt used in Spatial Transformation task w/ <summary> and <caption> tags in format.
Figure 13: Results of all methods on Qwen2VL-2B-Instruct, ID and DS performance at different training checkpoints.
Figure 14: Results of all methods on Qwen2VL-7B-Instruct, ID and DS performance at different training checkpoints.
Methods Steps Visual Counting
Clevr-Math (ID)
adversarial sub-multi addition subtraction AVG
Zero-Shot - 93.60 84.00 55.60 96.40 82.40
ANS-SFT 100 83.60 56.40 91.20 81.60 78.20
200 69.20 67.60 91.60 82.00 77.60
400 81.60 65.60 90.80 84.80 80.70
600 72.40 73.20 92.40 89.20 81.80
800 78.40 77.20 82.80 90.40 82.20
1200 85.60 78.00 91.60 95.60 87.70
1600 92.80 82.40 94.80 96.80 91.70
2187 95.20 92.80 97.60 99.20 96.20
CoT-SFT 100 49.20 40.00 82.00 69.20 60.10
200 65.20 55.60 88.00 76.40 71.30
400 66.00 57.20 90.00 79.60 73.20
600 67.20 59.20 87.20 82.80 74.10
800 77.60 61.60 92.40 85.20 79.20
1200 76.80 70.00 91.20 93.60 82.90
1600 80.80 66.80 91.60 92.00 82.80
2187 83.20 71.20 93.20 94.40 85.50
Reason-RFT-Zero 100 92.80 88.80 94.40 96.00 93.00
200 95.60 91.60 95.60 97.60 95.10
400 92.00 87.60 84.00 96.40 90.00
600 94.40 92.80 93.60 96.00 94.20
800 96.40 96.40 96.00 98.80 96.90
1200 98.40 95.60 100.00 99.60 98.40
1600 96.40 94.80 98.80 99.60 97.40
2500 98.40 95.60 99.60 100.00 98.40
Reason-RFT 100 89.60 73.20 93.60 95.60 88.00
200 89.20 78.00 95.20 96.40 89.70
400 92.80 82.40 95.20 97.60 92.00
600 94.80 86.00 96.80 97.20 93.70
800 96.80 88.40 96.80 98.80 95.20
1200 94.80 86.00 96.40 98.80 94.00
1600 94.40 91.60 97.20 99.60 95.70
2500 98.40 92.80 96.80 99.20 96.80
Table 10: Complete experimental results of Qwen2VL-2B-Instruct on the Clevr-Math test set after training on Clevr-Math. “sub-multi” donates the subtraction-multihop task.
Methods Steps Visual Counting
Super-Clevr-Math (DS)
addition subtraction add-sub sub-multi AVG DS-D DS-M
Zero-Shot - 10.40 54.40 0.00 63.20 32.00 42.67 0.00
ANS-SFT 100 51.20 37.60 11.60 39.20 34.90 42.67 11.60
200 38.40 55.60 8.40 15.60 29.50 36.53 8.40
400 40.80 45.20 5.60 35.20 31.70 40.40 5.60
600 41.20 61.60 8.00 35.60 36.60 46.13 8.00
800 49.20 50.40 7.20 26.00 33.20 41.87 7.20
1200 44.00 53.20 5.60 38.80 35.40 45.33 5.60
1600 48.80 53.60 6.00 26.00 33.60 42.80 6.00
2187 49.60 62.00 5.20 41.60 39.60 51.07 5.20
CoT-SFT 100 47.20 50.00 28.80 25.60 37.90 40.93 28.80
200 56.00 52.40 38.00 34.00 45.10 47.47 38.00
400 55.20 57.20 22.40 30.40 41.30 47.60 22.40
600 58.40 55.20 24.00 35.60 43.30 49.73 24.00
800 57.60 47.60 26.80 41.60 43.40 48.93 26.80
1200 58.00 54.40 35.60 32.40 45.10 48.27 35.60
1600 53.20 58.40 33.20 40.40 46.30 50.67 33.20
2187 53.60 58.80 36.80 36.80 46.50 49.73 36.80
Reason-RFT-Zero 100 46.00 65.20 6.80 58.80 44.20 56.67 6.80
200 48.80 66.00 9.20 57.60 45.40 57.47 9.20
400 42.00 71.20 8.40 50.80 43.10 54.67 8.40
600 47.20 65.20 7.60 47.60 41.90 53.33 7.60
800 56.40 69.20 6.80 55.20 46.90 60.27 6.80
1200 52.00 73.60 7.20 59.20 48.00 61.60 7.20
1600 51.60 71.60 6.40 54.80 46.10 59.33 6.40
2500 49.60 71.20 5.20 53.20 44.80 58.00 5.20
Reason-RFT 100 59.20 57.60 38.00 41.60 49.10 52.80 38.00
200 59.60 64.40 39.20 42.00 51.30 55.33 39.20
400 61.60 64.00 39.20 37.20 50.50 54.27 39.20
600 66.80 67.20 32.00 46.00 53.00 60.00 32.00
800 66.00 65.60 34.00 39.20 51.20 56.93 34.00
1200 67.20 65.20 33.60 40.80 51.70 57.73 33.60
1600 63.60 66.00 33.20 44.80 51.90 58.13 33.20
2500 68.00 67.20 28.40 44.80 52.10 60.00 28.40
Table 11: Complete experimental results of Qwen2VL-2B-Instruct on the Super-Clevr-Math test set after training on Clevr-Math. “add-sub” donates the addition-subtraction task, while “sub-multi” donates the subtraction-multihop task. “Direct Arithmetic”(DS-D) refers to the types of questions the model has previously seen during Clevr-Math training, while “Mixed Arithmetic”(DS-M) denotes the complicated types that the model has not encountered (i.e. questions with multi-step mixture of addition and subtraction).
Methods Steps Visual Counting
Clevr-Math (ID)
adversarial sub-multi addition subtraction AVG
Zero-Shot - 99.60 98.40 97.60 98.80 98.60
ANS-SFT 100 69.20 54.00 81.20 69.20 68.40
200 61.20 50.00 82.40 75.60 67.30
400 69.20 63.60 89.20 77.60 74.90
600 70.40 54.00 90.40 81.20 74.00
800 80.00 74.00 91.20 89.20 83.60
1200 86.80 79.20 94.40 91.20 87.90
1600 90.40 84.40 95.20 92.00 90.50
2187 96.80 89.20 96.80 97.20 95.00
CoT-SFT 100 81.60 63.60 91.20 83.60 80.00
200 80.00 64.00 92.00 88.80 81.20
400 72.40 66.00 88.80 79.60 76.70
600 77.60 66.00 94.40 89.20 81.80
800 78.40 65.20 94.00 87.20 81.20
1200 79.60 76.80 92.40 88.00 84.20
1600 86.40 78.00 92.80 93.20 87.60
2187 87.20 78.80 93.60 89.60 87.30
Reason-RFT-Zero 100 98.00 94.40 98.80 99.60 97.70
200 99.60 93.20 99.20 100.00 98.00
400 99.60 95.20 99.60 98.80 98.30
600 98.00 98.40 100.00 99.60 99.00
800 99.60 98.40 99.60 98.80 99.10
1200 100.00 98.00 99.60 99.20 99.20
1600 99.60 97.60 100.00 99.20 99.10
2500 99.60 98.40 100.00 99.60 99.40
Reason-RFT 100 88.80 79.20 95.60 94.40 89.50
200 92.00 80.00 96.40 95.20 90.90
400 94.40 84.40 96.00 95.60 92.60
600 92.80 84.00 96.40 97.60 92.70
800 92.80 85.20 96.80 96.40 92.80
1200 94.80 89.60 97.20 97.60 94.80
1600 94.80 86.40 97.60 97.20 94.00
2500 96.80 88.40 99.20 98.00 95.60
Table 12: Complete experimental results of Qwen2VL-7B-Instruct on the Clevr-Math test set after training on Clevr-Math. “sub-multi” donates the subtraction-multihop task.
Methods Steps Visual Counting
Super-Clevr-Math (DS)
addition subtraction add-sub sub-multi AVG DS-D DS-M
Zero-Shot - 46.80 75.20 4.80 41.60 42.10 54.53 4.80
ANS-SFT 100 57.60 41.20 5.60 46.40 37.70 48.40 5.60
200 42.00 38.80 8.00 33.60 30.60 38.13 8.00
400 37.20 46.40 5.20 31.60 30.10 38.40 5.20
600 32.00 44.80 12.40 19.20 27.10 32.00 12.40
800 38.80 38.00 6.80 37.20 30.20 38.00 6.80
1200 42.00 42.80 12.80 32.00 32.40 38.93 12.80
1600 36.40 48.40 11.20 17.20 28.30 34.00 11.20
2187 39.60 58.80 8.00 29.20 33.90 42.53 8.00
CoT-SFT 100 60.00 63.60 44.00 41.60 52.30 55.07 44.00
200 67.60 66.40 48.00 46.80 57.20 60.27 48.00
400 55.20 60.40 19.60 42.00 44.30 52.53 19.60
600 64.80 61.20 35.20 43.20 51.10 56.40 35.20
800 60.00 53.60 37.60 42.40 48.40 52.00 37.60
1200 51.20 56.00 35.20 39.60 45.50 48.93 35.20
1600 53.20 56.40 34.40 35.20 44.80 48.27 34.40
2187 51.60 51.60 33.60 32.80 42.40 45.33 33.60
Reason-RFT-Zero 100 58.80 82.80 24.00 62.40 57.00 68.00 24.00
200 56.00 83.20 18.80 50.00 52.00 63.07 18.80
400 62.40 79.60 22.80 37.60 50.60 59.87 22.80
600 61.20 85.20 17.20 49.20 53.20 65.20 17.20
800 52.80 86.80 20.40 52.00 53.00 63.87 20.40
1200 53.60 83.20 19.20 46.80 50.70 61.20 19.20
1600 61.20 84.80 18.40 43.20 51.90 63.07 18.40
2500 59.20 86.40 21.20 45.20 53.00 63.60 21.20
Reason-RFT 100 53.60 56.80 33.20 39.60 45.80 50.00 33.20
200 52.00 61.20 31.60 44.00 47.20 52.40 31.60
400 56.00 59.60 30.80 45.20 47.90 53.60 30.80
600 56.00 64.00 31.60 50.00 50.40 56.67 31.60
800 56.00 60.00 28.00 41.60 46.40 52.53 28.00
1200 66.00 65.60 38.00 50.40 55.00 60.67 38.00
1600 64.40 59.60 32.40 48.80 51.30 57.60 32.40
2500 62.80 60.80 35.60 44.80 51.00 56.13 35.60
Table 13: Complete experimental results of Qwen2VL-7B-Instruct on the Super-Clevr test set after training on Clevr-Math. “add-sub” donates the addition-subtraction task, while “sub-multi” donates the subtraction-multihop task. “Direct Arithmetic”(DS-D) refers to the types of questions the model has previously seen during Clevr-Math training, while “Mixed Arithmetic”(DS-M) denotes the complicated types that the model has not encountered (i.e. questions with multi-step mixture of addition and subtraction).
Methods Steps Structure Perception
Geometry3k (DS) GeoMath (ID)
CHOICE NON-CHOICE AVG CHOICE NON-CHOICE AVG
Zero-Shot - 40.25 1.00 20.63 35.57 20.31 25.86
ANS-SFT 100 35.25 16.25 25.75 58.72 29.89 40.37
200 33.25 17.50 25.38 56.38 35.44 43.05
400 30.75 17.00 23.88 64.77 35.06 45.86
600 - - - 73.83 38.12 51.10
800 32.75 16.00 24.38 72.15 36.40 49.39
1200 - - - 73.83 35.44 49.39
1600 29.00 16.00 22.50 74.83 37.36 50.98
1686 28.75 16.25 22.50 74.83 37.93 51.34
CoT-SFT 100 16.50 21.50 19.00 31.54 34.10 33.17
200 7.50 23.50 15.50 32.89 35.25 34.39
400 21.50 21.25 21.38 41.61 40.04 40.61
600 - - - 43.62 36.59 39.14
800 16.50 23.50 20.00 45.97 39.27 41.70
1200 - - - 53.02 40.04 44.76
1600 24.25 24.00 24.13 53.69 37.93 43.66
1686 26.75 23.75 25.25 51.34 38.31 43.05
Reason-RFT-Zero 100 32.25 17.75 25.00 41.61 31.23 35.00
200 33.00 18.50 25.75 48.99 35.06 40.12
400 41.50 23.50 32.50 52.68 34.87 41.34
600 37.00 22.75 29.88 60.74 37.55 45.98
800 42.25 25.00 33.63 62.42 40.42 48.42
1200 43.00 23.75 33.38 61.07 39.66 47.44
1600 42.75 22.25 32.50 63.09 38.31 47.32
1610 43.25 21.75 32.50 63.09 38.89 47.68
Reason-RFT 100 37.50 23.25 30.38 50.34 41.00 44.39
200 33.50 29.25 31.38 56.71 40.04 46.10
400 38.25 28.75 33.50 56.38 39.27 45.49
600 40.50 27.25 33.88 61.41 41.19 48.54
800 41.25 29.50 35.38 58.05 41.19 47.32
1200 40.25 31.00 35.63 61.74 42.34 49.39
1600 38.00 29.25 33.63 62.08 43.10 50.00
1610 36.75 29.50 33.13 60.74 42.34 49.03
Table 14: Complete experimental results of Qwen2VL-2B-Instruct on the Structure Perception task after training on GeoMath.
Methods Steps Structure Perception
Geometry3k (DS) GeoMath (ID)
CHOICE NON-CHOICE AVG CHOICE NON-CHOICE AVG
Zero-Shot - 45.25 23.00 34.13 61.07 38.12 46.46
ANS-SFT 100 38.50 18.25 28.38 64.77 34.87 45.74
200 32.50 22.75 27.63 69.46 35.25 47.68
400 - - - 72.48 40.42 52.07
600 32.25 18.00 25.13 73.49 39.27 51.71
800 - - - 75.50 37.93 51.58
1200 32.50 18.50 25.50 75.84 37.74 51.59
1600 32.50 18.25 25.38 75.84 37.36 51.34
1686 18.25 38.75 28.50 38.59 42.72 41.22
CoT-SFT 100 6.50 32.00 19.25 38.26 43.10 41.34
200 27.00 34.50 30.75 56.71 44.64 49.03
400 - - - 52.68 44.06 47.19
600 35.50 36.25 35.88 63.09 43.49 50.61
800 - - - 63.42 42.91 50.36
1200 29.50 37.50 33.50 64.09 44.06 51.34
1600 29.25 36.75 33.00 61.74 44.06 50.49
1686 58.50 41.75 50.13 56.71 45.98 49.88
Reason-RFT-Zero 100 59.00 44.25 51.63 63.42 45.21 51.83
200 62.00 43.00 52.50 70.47 45.40 54.51
400 - - - 70.13 46.74 55.24
600 64.75 45.25 55.00 70.47 49.23 56.95
800 - - - 66.11 46.17 53.42
1200 69.00 43.25 56.13 71.14 45.59 54.88
1600 66.25 43.25 54.75 69.80 46.55 55.00
1610 46.75 37.50 42.13 67.79 45.79 53.79
Reason-RFT 100 53.00 37.00 45.00 72.82 46.93 56.34
200 52.75 37.25 45.00 71.14 46.55 55.49
400 51.50 37.00 44.25 73.49 48.28 57.44
600 56.75 37.25 47.00 77.52 46.17 57.56
800 59.00 40.00 49.50 79.87 48.08 59.63
1200 56.00 39.50 47.75 74.50 49.62 58.66
1600 59.00 39.50 49.25 78.52 48.28 59.27
1610 59.00 39.50 49.25 78.52 48.28 59.27
Table 15: Complete experimental results of Qwen2VL-7B-Instruct on the Structure Perception task after training on GeoMath.
Method Steps Spatial Transformation
TRANCE (ID) TRANCE-L (DS-L) TRANCE-R (DS-R)
Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG
GPT-4o / 47.28 42.96 40.87 39.08 42.55 23.16 30.56 30.73 30.22 28.67 24.38 31.74 31.13 31.77 29.76
Zero-Shot / 2.10 3.27 4.08 5.68 3.78 2.02 4.73 5.57 6.08 4.60 2.39 4.73 5.59 5.95 4.67
ANS-SFT 100 15.90 19.12 14.67 13.12 15.70 10.60 11.33 10.01 9.07 10.25 11.08 12.17 10.17 10.07 10.87
200 23.97 29.56 33.98 33.95 30.37 13.25 26.76 31.26 32.58 25.96 13.53 26.54 30.53 31.20 25.45
400 44.95 42.58 40.75 33.65 40.48 26.03 35.98 34.26 29.38 31.41 24.06 35.62 35.46 31.73 31.72
800 62.10 56.55 53.01 47.55 54.80 24.70 42.27 43.05 42.65 38.17 26.09 38.98 42.95 42.55 37.64
1600 80.70 75.68 68.34 64.60 72.33 34.38 49.00 53.83 53.10 47.58 35.84 49.28 51.23 52.50 47.21
3200 82.85 80.30 78.00 71.60 78.19 36.22 52.61 55.23 54.27 49.58 38.51 52.78 54.47 53.98 49.94
final 82.70 79.93 76.70 70.22 77.39 36.00 52.82 54.59 53.55 49.24 39.63 53.75 54.33 53.60 50.33
COT-SFT 100 6.99 14.90 15.99 20.36 14.56 10.32 13.74 11.82 14.69 12.64 6.97 13.38 12.84 13.41 11.65
200 15.45 19.12 14.53 16.46 16.39 12.90 17.51 14.22 15.82 15.11 11.23 17.71 13.87 16.32 14.78
400 25.98 26.74 19.94 16.02 22.17 15.73 21.19 17.73 15.96 17.65 16.13 20.55 17.91 15.97 17.64
800 43.85 43.19 41.77 37.84 41.66 22.88 38.29 37.54 35.29 33.50 22.74 35.81 37.60 36.23 33.10
1600 52.82 61.06 54.38 45.85 53.53 28.65 43.95 40.99 40.13 38.43 29.00 41.23 40.78 39.26 37.57
3200 61.40 69.15 65.32 62.28 64.54 28.67 45.97 50.06 52.10 44.20 31.19 45.87 45.92 51.35 43.58
final 67.47 67.52 62.78 59.70 64.37 28.87 44.41 49.16 50.30 43.19 30.20 44.77 47.15 49.33 42.86
Reason-RFT-Zero 100 8.44 17.96 20.69 26.22 18.33 8.53 17.42 21.16 25.05 18.04 8.08 18.12 21.09 25.70 18.25
200 9.59 18.76 22.97 28.73 20.01 9.49 20.08 23.19 27.00 19.94 9.72 18.97 22.82 28.50 20.00
400 12.35 21.47 27.01 26.25 21.77 11.10 21.47 25.73 25.30 20.90 10.54 21.19 25.44 25.60 20.69
800 18.47 32.08 32.77 27.85 27.79 15.40 29.12 30.93 27.38 25.71 15.52 27.75 31.50 27.88 25.66
1600 36.78 40.20 37.78 34.51 37.32 19.96 33.03 35.84 34.49 30.83 20.39 32.85 33.87 33.90 30.25
3200 43.72 46.89 44.07 40.50 43.80 18.67 34.11 37.85 39.69 32.58 18.08 34.01 37.27 40.27 32.41
final 46.21 45.01 44.53 42.11 44.47 18.33 34.57 37.94 40.57 32.85 18.28 33.45 37.44 40.34 32.38
Reason-RFT 100 53.52 55.47 58.91 53.35 55.31 31.84 47.02 50.62 50.39 44.97 31.29 46.08 48.61 49.85 43.96
200 54.97 59.77 63.67 59.46 59.47 35.72 49.28 54.85 54.67 48.63 36.74 52.27 53.68 54.16 49.21
400 63.80 66.97 68.47 64.70 65.99 39.74 55.94 61.27 57.94 53.72 41.10 56.35 59.16 57.22 53.46
800 64.33 68.13 66.88 63.15 65.62 47.40 61.64 63.00 58.45 57.62 46.76 60.60 61.40 59.91 57.17
1600 76.47 73.42 74.05 69.16 73.28 52.68 62.22 66.56 64.71 61.54 53.47 65.17 64.98 63.52 61.79
3200 72.88 74.85 75.77 72.45 73.99 52.58 63.60 68.51 66.01 62.68 52.67 65.60 67.60 65.66 62.88
final 74.10 74.52 76.68 73.12 74.61 53.49 65.72 69.64 67.34 64.05 54.95 66.25 68.32 66.80 64.08
Table 16: Complete experimental results of Qwen2VL-2B-Instruct on the Spatial Transformation task after training on TRANCE.
Method Steps Spatial Transformation
TRANCE (ID) TRANCE-L (DS-L) TRANCE-R (DS-R)
Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG Level-1 Level-2 Level-3 Level-4 AVG
GPT-4o / 47.28 42.96 40.87 39.08 42.55 23.16 30.56 30.73 30.22 28.67 24.38 31.74 31.13 31.77 29.76
Zero-Shot / 16.25 16.42 10.96 10.48 13.53 11.71 16.80 11.50 10.85 12.72 13.30 16.08 10.55 11.18 12.78
ANS-SFT 100 40.30 37.05 30.67 28.35 34.09 32.07 31.12 26.00 26.80 29.00 26.38 29.71 27.74 26.48 27.58
200 65.18 53.33 49.43 45.15 53.27 33.29 45.14 45.61 45.52 42.39 35.17 43.43 45.97 43.02 41.90
400 65.33 59.35 57.17 50.77 58.16 32.40 44.13 47.69 46.23 42.61 32.10 44.88 47.04 45.25 42.32
800 78.90 70.67 63.97 62.10 68.91 34.08 50.62 51.99 52.95 47.41 34.22 50.40 50.62 52.88 47.03
1600 78.50 76.12 73.80 66.25 73.67 38.85 52.97 57.93 56.05 51.45 37.77 53.57 56.45 55.92 50.93
3200 83.80 83.23 82.83 78.17 82.01 40.10 56.02 61.02 59.90 54.26 40.78 55.06 61.67 60.98 54.62
final 83.70 84.10 82.50 78.45 82.19 39.67 55.58 61.84 60.05 54.29 42.64 54.84 61.44 60.38 54.83
COT-SFT 100 20.58 28.98 25.97 30.00 26.38 21.89 29.94 29.18 29.33 27.59 18.49 28.43 30.43 30.36 26.93
200 41.80 44.08 46.02 42.16 43.52 25.31 36.28 36.81 38.86 34.32 21.19 34.55 37.06 37.54 32.59
400 45.39 51.32 58.20 52.42 51.83 32.53 44.61 47.73 46.10 42.74 31.87 38.96 44.62 45.90 40.34
800 54.87 61.97 62.20 59.93 59.74 30.19 46.01 50.01 52.88 44.77 29.91 45.02 49.60 52.83 44.34
1600 71.27 71.14 72.82 69.93 71.29 28.82 46.43 51.01 58.94 46.30 29.08 45.25 52.24 58.31 46.22
3200 84.13 80.62 79.99 78.42 80.79 29.93 47.63 56.49 62.54 49.15 30.46 47.85 54.83 61.02 48.54
final 86.50 79.43 80.54 78.77 81.31 28.07 47.54 54.42 61.58 47.90 29.69 45.32 54.69 61.48 47.80
Reason-RFT-Zero 100 23.59 31.62 33.22 31.27 29.93 15.88 26.86 28.13 30.19 25.27 15.21 27.29 27.54 29.88 24.98
200 35.06 39.45 36.80 34.77 36.52 20.39 30.22 31.15 31.20 28.24 18.10 29.27 30.57 30.81 27.19
400 25.28 40.78 41.70 35.35 35.78 20.20 39.28 35.43 33.44 32.09 21.72 39.47 37.63 32.79 32.90
800 50.18 51.55 50.43 46.06 49.56 35.44 46.15 45.65 39.55 41.70 33.90 45.89 46.48 41.57 41.96
1600 59.60 61.90 57.30 55.36 58.54 43.95 55.03 52.96 50.60 50.64 41.28 56.46 51.64 49.08 49.62
3200 62.50 68.53 68.79 66.22 66.51 42.54 58.05 58.97 60.10 54.92 42.56 56.93 60.02 60.21 54.93
final 65.63 68.30 69.45 67.30 67.67 46.61 58.22 61.69 62.26 57.20 45.53 58.40 61.81 58.85 56.15
Reason-RFT 100 60.07 62.95 68.53 62.78 63.58 37.09 52.54 60.21 54.99 51.21 35.26 52.56 57.30 55.42 50.14
200 67.69 66.98 69.53 66.00 67.55 36.26 53.94 60.20 60.05 52.61 34.33 54.21 58.25 60.79 51.90
400 74.72 71.31 73.62 69.14 72.20 36.11 56.00 62.25 63.01 54.34 36.95 55.72 60.82 64.88 54.59
800 72.57 71.04 75.58 73.26 73.11 38.93 56.86 64.10 63.69 55.90 35.06 54.39 60.82 64.88 53.79
1600 79.18 76.70 76.59 73.26 76.43 37.25 53.44 66.38 66.76 55.96 35.01 55.39 66.02 66.67 55.77
3200 79.70 79.22 80.13 76.61 78.92 37.63 55.73 65.24 68.02 56.66 37.18 56.34 63.90 68.89 56.58
final 80.48 79.15 81.97 78.26 79.97 43.68 61.04 64.66 68.05 59.36 45.45 56.34 63.33 69.33 58.61
Table 17: Complete experimental results of Qwen2VL-7B-Instruct on the Spatial Transformation task after training on TRANCE.
Figure 15: Visualization examples of Reason-RFT and CoT-SFT on general visual reasoning.
Figure 16: Visualization examples of Reason-RFT and CoT-SFT on general visual reasoning.
Figure 17: Visualization examples of Reason-RFT and CoT-SFT on general visual reasoning.
Figure 18: Visualization examples of Reason-RFT and CoT-SFT on general visual reasoning.
Figure 19: Visualization examples of Reason-RFT and CoT-SFT on Visual Counting task.
Figure 20: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 21: Visualization examples of Reason-RFT and CoT-SFT on Structure Perception task.
Figure 22: Visualization examples of Reason-RFT and CoT-SFT on Spatial Transformation task.
Figure 23: Visualization examples of Reason-RFT and CoT-SFT on Spatial Transformation task.